### v1

# TrajAgent: An Agent Framework for Unified Trajectory Modelling

Abstract.
Trajectory modeling, which includes research on trajectory data pattern mining and future prediction, has widespread applications in areas such as life services, urban transportation, and public administration. Numerous methods have been proposed to address specific problems within trajectory modelling. However, due to the heterogeneity of data and the diversity of trajectory tasks, achieving unified trajectory modelling remains an important yet challenging task. In this paper, we propose TrajAgent, a large language model-based agentic framework, to unify various trajectory modelling tasks. In TrajAgent, we first develop UniEnv, an execution environment with a unified data and model interface, to support the execution and training of various models. Building on UniEnv, we introduce TAgent, an agentic workflow designed for automatic trajectory modelling across various trajectory tasks. Specifically, we design AutOpt, a systematic optimization module within TAgent, to further improve the performance of the integrated model. With diverse trajectory tasks input in natural language, TrajAgent automatically generates competitive results via training and executing appropriate models. Extensive experiments on four tasks using four real-world datasets demonstrate the effectiveness of TrajAgent in unified trajectory modelling, achieving an average performance improvement of 15.43% over baseline methods.
*Both authors contributed equally to this research.
copyright: acmcopyrightcopyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
With the rapid development of web services and mobile devices (Zheng, 2015; Chen et al., 2024), large-scale trajectory data, such as check-in data from social network (Yang et al., 2016), have been collected, greatly facilitating research in trajectory modelling. Trajectory modelling involves the processing, mining and prediction of trajectory data, with widespread applications in urban transportation, location services and public management. The typical areas of trajectory modelling (Chen et al., 2024; Graser et al., 2024) can be classified into three main categories: trajectory classification (Liang et al., 2022), trajectory prediction (Yang et al., 2022), and trajectory generation (Wei et al., 2024). Each category encompasses various sub-tasks; for instance, the trajectory prediction task can be further divided into next location prediction task (Liu et al., 2016), final destination prediction task (Zhang et al., 2019), and travel time estimation task (Wang et al., 2018), among others. Given the huge value of trajectory modelling in diverse practical applications, various algorithms and models (Graser et al., 2024) have been proposed to address these tasks, particularly deep learning-based models in recent years. This has facilitated significant advancements in the field, with many tasks achieving a high level of performance.
However, existing methods are designed for specific tasks and datasets, making it difficult to share them across different tasks and data sources. For example, TrajFormer (Liang et al., 2022) is tailored for trajectory classification and cannot be applied in trajectory prediction or trajectory generation. Flash-back (Yang et al., 2020) is designed for sparse check-in trajectory prediction and is not suitable for dense GPS trajectory or road network-based trajectory modelling. In other words, due to the heterogeneity of application scenario and the diverse nature of trajectory data–varying in resolution, format and geographical regions–existing methods can only be applied in limited task with specific data for specific regions. Without a unified model, the deployment and application of these advanced models become costly and inefficient, hindering further development of trajectory modelling area. Therefore, developing a unified model for various trajectory modelling tasks is both emergent and fraught with significant challenges.
In last two years, the rapid development of large language models (LLMs) (OpenAI, 2024; Touvron et al., 2023) with extensive commonsense and powerful reasoning abilities presents great potential for solving challenging tasks, such as mathematics (Romera-Paredes et al., 2024) and logic games (Yao et al., 2022b). Furthermore, LLM based agent (Xi et al., 2023; Sumers et al., 2023) have emerged as a new paradigm, offering more robust and intelligent task-solving capabilities. These agents have demonstrated stable and promising performance in tackling complex real-world tasks, such as web-navigation (Yao et al., 2022a; Nakano et al., 2021; Zhou et al., 2023) and software development (Hong et al., 2023; Qian et al., 2024; Yang et al., 2024). Leveraging the powerful understanding and reasoning abilities of LLM based agents, researchers explore the potential of utilizing LLM based agent for the unified modelling of diverse machine learning tasks (Huang et al., 2024; Shen et al., 2024; Zhang et al., 2024; Wu et al., 2024). For example, HuggingGPT (Shen et al., 2024) utilize LLM as a core manager to address various machine learning tasks with existing AI models, VisionLLM (Wu et al., 2024) investigates unified modelling for vision tasks across different domains, while Mora (Yuan et al., 2024) designs a multi-agent framework for replicating Sora’s functionalities with combining existing open source models.
Inspired by these works, we explore the potential of addressing the unified trajectory modeling problem using an LLM-based agent framework. However, designing an LLM-based agent for unified trajectory modelling is challenging due to several factors. Firstly, trajectory data and existing models for various trajectory modelling tasks are diverse and often inconsistent, making the integration and deployment of these models difficult. Secondly, the processing steps from raw trajectory data to final output results are lengthy and cumbersome (Chen et al., 2024), resulting in a large action space for the planning and execution of entire process. Third, the performance of existing models highly depends one data adaptation and parameter optimization, making the direct usage of existing pre-trained models infeasible.
Figure 1. Framework of TrajAgent, including UniEnv with unified data and model interface for running experiment, TAgent an agentic workflow for executing the whole trajectory modeling procedure, and AutOpt for the model enhancement and adaptation on selected data.
In this paper, we propose TrajAgent, a systematic agent framework for unified trajectory modelling and analytics. Within TrajAgent, we first design a unified environment, UniEnv, for processing diverse trajectory data and preparing a cohesive running environment for various trajectory modelling tasks. In UniEnv, we define unified data and model interfaces to support the easy execution of various methods for trajectory modelling. Next, we develop an agentic workflow called TAgent for the automatic planning and execution of trajectory modelling tasks. In TAgent, we decompose the task workflow into four steps: task understanding, task planning, task execution and task summarization. In each step, we design a specialized expert agent to perform specific operations. Notably, we introduce AutOpt to enhance the task execution module, enabling the adaption of model performance through various automatic optimization mechanisms, including data augmentation, parameter optimization and joint optimization. In summary, our contributions are as follows,
  • To the best of our knowledge, TrajAgent is the first attempt to employ an LLM-based agentic framework for unified trajectory modeling across diverse trajectory data and various trajectory tasks.
  • In TrajAgent, we design a agentic workflow to decompose the whole trajectory modelling task into several sub-tasks, all carried out within a unified environment UniEnv. Finally, we develop a comprehensive model optimization module to further enhance the model performance on targeted data.
  • Extensive experiments on four representative trajectory modelling tasks across four real-world datasets demonstrate the effectiveness of proposed framework, which the optimized model achieving an average performance gain of 15.43%.
2. Methods
Figure 1 presents the whole framework of TrajAgent. It contains three components: 1) UniEnv, an environment with a unified data and model interface that supports the execution and training of various models; 2) TAgent, an agentic workflow designed to automatically complete diverse trajectory modelling task. It accepts requirements in the form of natural language input and directly outputs the adapted models; 3) AutOpt, an additional optimization module for TAgent that enhances the performance of specific model through data and parameter adaptation. Details of each component are presented as follows.
2.1. UniEnv
As shown in the left of Figure 1, UniEnv is a comprehensive and integrated environment that bridges trajectory data, tasks, and models, providing a foundational platform for trajectory modelling and analysis. It is designed to support the entire lifecycle of trajectory modelling workflows, from data preparation to task execution and model optimization. UniEnv comprises four key components: a rich set of datasets accompanied by processing tools, a comprehensive task collection that defines and manages various task types, a extensive model library with available source code, and an external tools pool for extending the capabilities of TrajAgent. Each component is seamlessly connected through a unified interface, enabling agents to plan and execute trajectory modeling tasks with minimal complexity.
2.1.1. Task Interface
We summarize the supported trajectory modelling tasks in UniEnv as follows.
  • Trajectory prediction task: next location prediction and travel time estimation.
  • Trajectory classification task: trajectory user linking and trajectory anomaly detection.
  • Trajectory generation task: trajectory recovery and trajectory generation.
To enhance the clarity and effectiveness of understanding users’ language queries, we provide a detailed language description for each task. This description helps extract precise task requirements from user queries and facilitates subsequent data and model selection processes.
2.1.2. Data Interface
UniEnv supports two commonly used trajectory data formats, namely Checkin trajectory (i.e., sequence of visited POIs) and GPS trajectory (i.e., sequence of gps points). These datasets, which come from different cities and with distinct forms, are pre-processing through a standard pipeline that ensures compatibility across the system. Pre-processing steps are done by generated code scripts from LLMs, include data cleaning, normalization, format transformation, which are crucial for handling inconsistencies between real-world datasets and task models. After processing, we will add a description for each dataset to support efficient data selection in the subsequent stage.
2.1.3. Model Interface
To support various model training in TrajAgent, we select at least one well-known model for each task and adapt them to match the running environments in UniEnv. Furthermore, we extract the semantic context from original paper of each model with the txyz.ai APIs  [^1] https://www.txyz.ai/ to generate detailed description of model. In this description, the verified data information and supported trajectory task information are recorded which supports the data and model selection in the subsequent trajectory planning stage of TrajAgent.
2.1.4. External Tools
To extend the capabilities of TrajAgent, we collect several external tools in UniEnv, including paper context extraction tool txyz.ai, hyperparameter optimization tool optuna  [^2] https://optuna.org/, processing and visualization tool for trajectory data movingpandas  [^3] https://github.com/movingpandas/movingpandas, open street map data processing tool osmnx  [^4] https://github.com/gboeing/osmnx.
Here, we also regard the LLM APIs utilized in the agentic workflow as one of the interface in the UniEnv, including the ChatGPT API  [^5] https://openai.com/ and DeepInfra  [^6] https://deepinfra.com for open-source LLMs.
2.2. TAgent
As shown in the middle of Figure 1, the architecture of TAgent is organized into four key modules: Task Understanding, Task Planning, Task Optimization, and Task Summary, which form an automated processing chain from user query to final result, eliminating the need for human-in-the-loop. Specifically, the task understanding module first receives user instructions in natural language form, and analyzes and identifies the type, name, and other key information of the tasks involved. Then, the task planning module will plan for the identified task, including dataset matching and model selection. Next, the task execution module executes the planning task and cooperate with the additional model optimization module AutOpt to further improve the task performance from both data augmentation and parameter selection perspectives. Last, the task summary module generate an analytical report of the task based on historical interactions and decisions of TAgent. Following the common practice(Shinn et al., 2024), each module in TAgent can be regarded as a small agent, which consists of a function module for executing its core function, a memory for recording the history interaction, and a reflection module for learning practice from the memory.
2.2.1. Task Understanding Module
As the first module of TAgent, task understanding module is designed to interact with user and extract detailed task information to launch subsequent stages. Given the user query, understanding module recognize the potential task name from it with the predefined supported tasks as additional input. If users ask for the out-of-scope tasks which has not been supported in the UniEnv, we will directly recommend user to select task from the supported list.
2.2.2. Task Planning Module
Follow the task understanding module is the planning module which is designed to generate the subsequent execution plan for efficient experiments of trajectory modelling. The input of the planning module is the task name and description from the undertanding module, the supported data and model with brief description from UniEnv. With the carefully designed prompt, the generated execution plan will contain the data name and model name for the given task, and also the detailed model optimization plan.
Due to the characteristics of different tasks and existing practice, not all the model optimization are necessary for each task. If possible, skipping the optimization step which is time-consuming and costly can accelerate the whole procedure without sacrificing performance. For example, while the data augmentation is very important for sparse check-in trajectory data, it is not so useful for dense GPS trajectory in many applications. Thus, we can generate simplified execution plan for dense GPS trajectory in some tasks to improve the success rate of the whole task and also reduce the cost. After generating the plan, it will start a simple execution step to verify the feasibility of the plan. Once any error occurs during the execution, e.g., the model name is wrong, the planning module will obtain the feedback from UniEnv and start to regenerate a new plan with the last plan as the failed history in its memory.
2.2.3. Task Execution Module
Give the execution plan, the task execution module is responsible for invoking UniEnv to execute the experiment plan. In addition to the previously mentioned basic execution interface, another interface of this module is to call AutOpt module to complete the model optimization according to the optimization plan. For both interface, task execution module will give the feedback including error information for failed cases and performance metrics for success cases. Detailed designs of AutOpt are introduced in the following section 2.3.
2.2.4. Task Summary Module
After the execution module, we design a task summary module to analyze the execution records to generate the optimization summary of the task. The summary contains the optimization path during the experiment and the final optimization result for the given task. User can also directly utilize the optimized model from the experiments via APIs for further applications.
2.3. AutOpt
Due to the geospatial heterogeneity and the diversity of trajectory data, the trajectory models usually cannot be directly transferred between data and regions. In other words, for various data in different region, the model needs to be trained from scratch. Thus, the sufficient optimization of various models becomes emergent. In TrajAgent, we design AutOpt module to complete this task. As shown in the right part of Figure 1, AutOpt provides several optimization mechanisms: data augmentation, parameter optimization, joint optimization, prompt optimization and optimization via tools. Following Reflexion (Shinn et al., 2024), we design optim agent for each kind of optimization mechanism. The standard optim agent utilizes the history operation and related results as the feedback to update its action in the next step. Specifically, it works as two stages, including ”think” and ”action”. To support the ”think then action”, it builds a long-term memory for recording all experimental data and a short-term memory for historical actions. In the ”think” stage, optim agent analyzes the long-term memory and meta information of experiment and generating the guidance of action in the short-term memory. In the ”action” stage, optim agent analyze the results in short-term memory to generate the action. Different optimization mechanism utilize the same optim agent with different action space and optimization tips.
2.3.1. Data Augmentation
Based on the standard optim agent, we introduce the specific action space for data augmentation. Following practice from existing works (Zhou et al., 2024; Dang et al., 2023; Yin et al., 2024), we define a fixed set with ten operators for data augmentation, e.g., insert, replace, split and so on. During the optimization, the operator set with simple description is provided to the optim agent, it can select any combination of operators with their simple parameters as the action. After action, optim agent obtain the feedback information, e.g., performance metrics, from UniEnv to continue update its memory and action.
2.3.2. Parameter Optimization
The action space of parameter optimization is defined based on the parameters of model itself. We define a parameter configuration file for each model, the optim agent reads the configuration file and generates code as the action to update the parameters in it. To better understand the meaning of each parameter, we add comments for each parameter in the file. While the action space is large, this kind of action space is flexible to adapt with any models.
2.3.3. Joint Optimization
Furthermore, we introduce the joint optimization mechanisms in AutOpt to improve the final performance. Due to the different working paradigms, the direct combination of two kinds of optimizations is unsuccessful. We designate the optimization order to prioritize data augmentation first, followed by parameter optimization. This means that once the performance of data augmentation stabilizes, the agent proceeds with parameter optimization. This procedure can be repeated a fixed number of times until it meets the stop criteria.
3. Experiments
In this section, we first introduce the experimental settings, including the datasets, task models, and evaluation metrics. Next, we present the experimental results, covering the performance across multiple trajectory tasks as well as the effectiveness of agent execution. Following that, we perform an ablation study to assess the impact of different components within the proposed framework. Subsequently, we investigate the effect of key parameters for TrajAgent execution. Finally, we conclude with a case study to further illustrate the practical application of our approach.
3.1. Settings
3.1.1. Dataset
  • Foursquare (FSQ): This dataset consists of 227,428 check-ins collected in New York City from 04/12/2012 to 02/16/2013, Each check-in is associated with a timestamp, GPS coordinates and corresponding venue-category.
  • Brightkite (BGK): This dataset contains 4,491,143 check-ins of 58,228 users collected from BrightKite website.
  • Porto: This dataset contains 1.7 million taxi trajectories of 442 taxis running in Porto, Portugal from 01/07/2013 to 30/06/2014. Each trajectory corresponds to one completed trip record, with fields such as taxiID, timestamp and the sequence of GPS coordinates.
  • Chengdu: This dataset contains GPS trajectory records of Chengdu from 01/11/2016 to 30/11/2016. Each record includes taxiID, timestamp, longitude and latitude, collected and released by Didi Chuxing.
  • UserQueries: To verify the effectiveness of the whole system, we utilize self-instruct method (Wang et al., 2022) with 5 seed queries to generate 300 user queries as the experiment input.
More detailed information for the Check-in and GPS trajectory datasets are summarized in Table 1 and Table 2, respectively. Note that the raw datasets are typically city-scale and data-intensive. Loading them all into the TrajAgent framework is costly. In this work, we are selecting a part of data for task model training and testing during experiments.
Table 1. Statistics of the Check-in trajectory datasets.
Datasets Foursquare (FSQ) Brightkite(BGK)
Num. Users 463 272
Num. POIs 19870 50061
Num. Trajectories 10632 22208
Table 2. Statistics of the GPS trajectory datasets.
Datasets Porto Chengdu
Sampling Rate 15s 3s
Num. Trajectories 1,710,670 5,819,383
Avg. Trajectory Length (m) 3522.64 2857.81
Avg. Travel Time (s) 724.20 436.12
Latitude Range [41.1401, 41.1859] [30.6529, 30.7277]
Longitude Range [-8.6902, -8.5491] [104.042, 104.129]
3.1.2. Task models
In this work, We adopt various AI models and incorporate them into TrajAgent for solving trajectory-related tasks. According to the type of tasks they deal with, these models can be framed in the following categories:
  • Models for Trajectory Prediction. We consider the classic method RNN (Liu et al., 2016) and well-performing approaches FPMC (Rendle et al., 2010) and DeepMove (Feng et al., 2018) for the task of next location prediction.
  • Models for Trajectory User Linkage. We focus on methods for solving the task of Trajectory User Linkage under check-in trajectories, such as MainTUL (Chen et al., 2022) and DPLink (Feng et al., 2019).
  • Models for Travel Time Estimation. We consider the mainstream model DeepTTE (Wang et al., 2018) to estimate the travel time for a given GPS trajectory.
  • Models for Trajectory Anomaly Detection. We considered the well-performing method GMVSAE (Liu et al., 2020) for detecting anomalous trajectory.
3.1.3. Metrics
To evaluate the performance of all models on multiple trajectory tasks, we employ the following different metrics:
  • Acc@5 and Hit@5: Acc@5 refers to the percentage of the first five results predicted correctly. Hit@5 measures whether at least one of the top-5 predictive results is correct.
  • MAE and AUC: Mean Absolute Error (MAE) indicates the amount of deviation from the actual values. Area Under ROC Curve (AUC) measures how well the model correctly distinguishes the type of the sample.
In our experiments, Acc@5 is used to measure the accuracy of trajectory prediction and agent execution, Hit@5 is adopted for evaluating the performance of trajectory user linkage task; MAE is employed to compute the error of travel time estimation, while the metric AUC is used to assess the performance of trajectory anomaly detection.
3.2. Main Results
In this section, we evaluate the effectiveness of TrajAgent from two dimensions: 1) Performance of trajectory modelling tasks accomplished through TrajAgent, and 2) Execution efficiency of each stages in TrajAgent.
3.2.1. Performance on Trajectory-related Tasks
Here, we focus on the performances of proposed TrajAgent framework across four types of trajectory tasks: Trajectory Prediction (TP), Trajectory User Linkage (TUL), Travel Time Estimation (TTE) and Trajectory Anomaly Detection (TAD). The first two tasks, TP and TUL, are based on check-in trajectory data, while TTE and TAD utilize GPS trajectory data. The results for these tasks are summarized in Table 3 and Table 4, respectively.
For each task, we report the model performance under four different configurations: the original model (Origin), with Data Augmentation (+DA), with Parameter Optimization (+PO), and with joint optimization (+JO). This setup allows us to assess the individual and combined effects of data augmentation and parameter optimization on the task performance.
Table 3 presents the performance comparison on check-in trajectory tasks TP and TUL. In the TP task on FSQ dataset, DeepMove improves from 0.3422 (Origin) to 0.4018 (+JO), a 17.42% increase. RNN, as a simpler model, experiences the most significant gains, with accuracy increasing by 51.36% through joint optimization. On the BGK dataset, DeepMove continues to perform well, achieving 0.6100 (+JO), marking a 9.52% improvement. Across both datasets, TrajAgent’s optimization techniques consistently enhance model performance, with the FSQ dataset showing particularly notable improvements. For the TUL task, MainTUL and DPLink exhibit consistent improvements under TrajAgent’s optimization strategies. the performance of MainTUL increases from 0.4871 to 0.6121 (+JO) in FSQ, while DPLink achieves a more moderate increase of 6.08%. These results demonstrate the effectiveness of the TrajAgent framework, particularly in automating check-in trajectory tasks. The results also highlight the utility of TrajAgent in terms of integrating data enhancement and parameter optimization strategies for better problem solving.
Table 3. Comparison of task performance on check-in trajectories for TrajAgent with different configurations. ‘DA’ represents Data Augmentation, ‘PO’ denotes Parameter Optimization, ‘JO’ indicates Joint Optimization. $\delta$ represents improvements.
Task TP TUL
Model RNN FPMC DeepMove MainTUL DPLink
Metrics Acc@5 Hit@5
FSQ Origin 0.1795 0.2939 0.3422 0.4871 0.7551
+DA 0.2667 0.2939 0.4018 0.5973 0.7551
+PO 0.1795 0.3066 0.3422 0.5691 0.7686
+JO 0.2717 0.3066 0.4018 0.6121 0.8010
$\delta$ 51.36% 4.32% 17.42% 25.66% 6.08%
BGK Origin 0.4422 0.5241 0.5570 0.5908 0.8993
+DA 0.5416 0.5275 0.5647 0.6836 0.9613
+PO 0.5022 0.5250 0.6041 0.6683 0.9552
+JO 0.5470 0.5275 0.6100 0.7145 0.9622
$\delta$ 23.70% 0.65% 9.52% 20.94% 6.99%
Table 4 presents the performance comparison on GPS trajectory tasks, specifically focusing on TTE and TAD tasks. The models evaluated are DeepTTE for TTE and GMVSAE for TAD, with results shown for the original configuration (Origin) and after Joint Optimization (+JO). For the Proto dataset, the original model has an MAE of 8.48, which significantly improves to 5.85 after joint optimization, resulting in a 31% reduction in prediction error. On the Chengdu dataset, the MAE decreases from 7.23 to 5.95. As for TAD task, we find that the AUC scores were already high, the small but consistent improvements in both datasets suggest that TrajAgent’s joint optimization can further refine model performance, even for tasks where models initially perform well.
Table 4. Comparison of task performance on GPS trajectories for TrajAgent with different configurations.
Task
TTE
TAD
Model DeepTTE GMVSAE
Metrics MAE AUC
Porto origin 8.48 0.9892
+JO 5.85 0.9899
$\delta$ 31.01% minor
Chendu Origin 7.23 0.978
+JO 5.95 0.984
$\delta$ 17.70% 0.61%
3.2.2. Efficiency of Agentic Workflow
We select the trajectory prediction task as an example to demonstrate the efficiency of agentic workflow.
Table 5. Execution efficiency of TrajAgent implemented by different LLMs.
LLM Extraction Processing Data/Model Selection Data Augmentation. Parameter Optim. Joint Optim.
Succ. Succ. Succ. Acc@5 Succ. Acc@5 Succ. Acc@5 Succ. Acc@5
Qwen2-7B 85.00% 30% 72% 83.33% 15% 0.2015 25% 0.1833 64% 0.2668
Mistral7B-V3 78.89% 42% 88% 90.91% 94% 0.2940 95% 0.2087 82% 0.2980
LLama3-8B 69.44% 28% 81% 80.25% 18% 0.1790 11% 0.1809 65% 0.2809
Gemma2-9B 83.88% 12% 57% 52.63% 18% 0.1822 15% 0.1848 70% 0.2970
Gemma-2-27B 79.44% 30% 70% 88.57% 78% 0.2507 30% 0.1775 78% 0.3366
GPT3.5-Turbo 88.89% 54% 100% 82.00% 88% 0.2846 90% 0.1809 92% 0.3295
LLama3-70B 83.33% 100% 95% 86.32% 92% 0.2931 83% 0.1848 95% 0.3473
Qwen2-72B 92.22% 95% 100% 95% 96% 0.3925 70% 0.1816 94% 0.4333
GPT-4o-mini 95.56% 92% 100% 98.00% 90% 0.2967 85% 0.1822 96% 0.3724
Table 5 compares the execution efficiency of TrajAgent implemented by different LLMs, across each stages in the trajectory modelling workflow. We can observe that Qwen2-72B and GPT-4o-mini demonstrate the highest success rates (i.e., Succ.) across key stages such as Data Extraction, Processing, and Data/Model Selection, with over 90% success in each. They also excel in Joint Optimization, with success rates of 94% and 96%, resulting in superior Acc@5 scores of 0.4333 and 0.3724 respectively. In contrast, models like Gemma-2-9B and LLama3-8B struggle with lower processing and data selection success rates, which results in reduced overall performance. Their weaker performance in key optimization stages, especially in parameter selection, reflects their limitations in effectively supporting TrajAgent.
These results show that TrajAgent’s efficiency is strongly influenced by the base LLM, with high-performing models like Qwen-72B and GPT-4o-mini significantly enhancing its capabilities. To fully realize TrajAgent’s potential, it is essential to improve LLM performance in data augmentation and parameter optimization, as many LLMs currently struggle in these aspects.
3.3. Ablation Study
Table 6. Ablation Study of TrajAgent. ‘MS’ stands for Model Selection, ‘DA’ represents Data Augmentation, ‘PO’ denotes Parameter Optimization, ‘JO’ stands for Joint Optimization. $\downarrow$ indicates a decrease in performance, and $\uparrow$ indicates an improvement.
Agent Variants MS DA PO JO
Succ. Acc Succ. Acc Succ. Acc Succ. Acc
TrajAgent 100% 98% 98% 0.3050 89% 0.1895 85% 0.3233
w/o Reflection 100% 95$\downarrow$ 98% 0.3028$\downarrow$ 90$\uparrow$ 0.1872$\downarrow$ 82$\downarrow$ 0.3212$\downarrow$
w/o Memory 99$\downarrow$ 80$\downarrow$ 85$\downarrow$ 0.1707$\downarrow$ 70$\downarrow$ 0.2050$\uparrow$ 68$\downarrow$ 0.1804$\downarrow$
To assess the effectiveness of the key components in TrajAgent, we conducted an ablation study using execution efficiency as the evaluation criterion. Specifically, we examined the framework by isolating two main designs: the memory unit and the reflection mechanism, resulting in two variants: 1) w/o Reflection, where the reflection mechanism is removed, and 2) w/o Memory, where the memory unit is excluded. Additionally, it is important to note that GPT-4o-mini was used as the base LLM for all variants of TrajAgent in this study. The experimental results are presented in Table 6. We can find that: 1) Removing either component leads to average performance declines, with the memory unit being especially critical for maintaining execution efficiency. For example, removing the memory unit causes significant drops in accuracy for both Data Augmentation (DA) and Parameter Optimization (PO), underscoring its essential role in these processes. 2) Interestingly, removing a component occasionally results in slight increases in success or accuracy, suggesting that some components may introduce overhead or complexity in specific stages. Overall, the combined use of the memory and reflection mechanisms is crucial for optimizing TrajAgent’s performance.
3.4. Parameter Study
In this section, we analyze the effects of two important parameters in the TrajAgent framework: 1) thought step: the number of steps that agent thought before taking action, and 2) memory size: the size of memory units in TrajAgent. To be specific, we vary the values of these parameters to investigate their impact on agent performance, and the results are illustrated in Figure LABEL:fig:param-test(a) and (b). As shown in Figure LABEL:fig:param-test(a), increasing thought steps significantly enhances Joint Optimization (JO) and Data Augmentation (DA). This indicates that giving TrajAgent more reasoning time improves its overall performance. However, the peak performance for Parameter Optimization (PO) and JO at around 10 steps suggests that once sufficient contextual understanding is achieved, further increasing the thought steps provides diminishing returns and may not be beneficial. From Figure LABEL:fig:param-test(b), we find that increasing memory size initially improves performance for JO and DA, with peak accuracy reached at a memory size of 10. This suggests that having sufficient memory allows TrajAgent to store and utilize relevant information effectively. However, beyond this point, performance starts to decline, indicating that excessive memory may introduce unnecessary complexity or irrelevant data, leading to diminished returns for JO and DA. For Parameter Optimization (PO), memory size has little effect, reflecting that this task does not rely as heavily on memory records.
3.5. Case Study
In this section, we analyze some cases where the TrajAgent’s optimization performance is suboptimal. The overall process is illustrated in Figure 3. The optimization module is a key component affecting overall performance.
The first issue is optimization trap present in data augmentation module of TrajAgent. Specifically, it refers to the situation where agent sometimes ignores the contents of the memory during the thought process. Even when the chosen parameter combination yields poor training results, the model overlooks the error feedback (i.e., the ”Not good enough…” in the memory). The ”optimization trap” occurs even in the best-performing GPT-4o-mini. As the length of the Memory increases, the impact of the optimization trap on overall accuracy grows. Once an optimization trap occurs, all memories within the same step tend to favor the same ineffective combination. We believe the causes of the optimization trap could be: (1) excessively long prompts leading to truncated inputs; (2) insufficient proportion of memory in the total input.
The second issue is the sub-optimality appears in parameter optimization module of TrajAgent. This phenomenon exists across various datasets and model sizes. We believe the reasons for the poor performance might be: (1) the TrajAgent parameter optimization module is overly sensitive to the format of model outputs, treating all responses that do not meet the format requirements as invalid; (2) adjustments to certain parameters result in increased training time, reducing the total number of iterations.
Figure 3. A representative example case of TrajAgent.
4. Related Work
4.1. Trajectory Modelling and Analytics
In recent year, trajectory modelling (Zheng, 2015; Chen et al., 2024; Graser et al., 2024) makes great progress on its core research questions, including prediction (Liu et al., 2016; Feng et al., 2018; Shen et al., 2020), classification (Liang et al., 2022; Song et al., 2018), and generation (Yuan et al., 2022).
Trajectory prediction is to predict the next location or the trip destination of the trajectory. ST-RNN (Liu et al., 2016) is the first work for applying recurrent neural network into the trajectory prediction task. Following this work, lots of works have been done to solve the critical issues of trajectory prediction with deep learning, for example, DeepMove (Feng et al., 2018) for multi-granularity regularity modelling, Flash-back (Yang et al., 2020) for sparse trajectory prediction, GETNext (Yang et al., 2022) for utilizing aggregated mobility pattern. Besides, to solve the travel time prediction task, another important prediction problem for trajectory modelling, deep learning models including DeepTTE (Wang et al., 2018), TTPNet (Shen et al., 2020) and CARTE (Huang et al., 2022) are proposed and achieve promising performance in various scenarios. As for the classification problem in trajectory modelling, the most well-known tasks are trajectory mode detection (Jiang et al., 2017) and trajectory user linkage (Gao et al., 2017) task. Various techniques, such as siamese network (Feng et al., 2019), distillation learning (Chen et al., 2022) and transformers (Liang et al., 2022) are successfully applied to discover the pattern among trajectories to solve the classification tasks. Finally, due to the natural limitation of trajectory data quality, trajectory data recovery (Xi et al., 2019; Xia et al., 2021) and generation (Ouyang et al., 2018) are also important tasks in trajectory modelling which are crucial for the practical application in the real world. Thus, follow the the trend of generative modelling, diverse generative models including GAN (Ouyang et al., 2018), diffusion models (Yuan et al., 2022; Wei et al., 2024) are utilized to improve the modelling of human dynamics.
While these specific methods accelerate the development of trajectory modelling from different aspects, they can only handle one type of task. In other words, the unified model for all the trajectory modelling task is still missing due to the heterogeneity of tasks and trajectory data. In this paper, we propose to utilize the power of LLM and agent to build a unified model framework for diverse trajectory modelling tasks, which can automatically handle various trajectory data and trajectory modelling tasks without human intervention.
4.2. Large Language Model and Agent
With the advent of ChatGPT (OpenAI, 2024), large language models(LLMs) with extensive commonsense and outstanding reasoning abilities have been widely explored in many domains, such as mathematics (Yao et al., 2022b), question answering (Zhuang et al., 2023), and human-machine interaction (Touvron et al., 2023). Furthermore, agentic framework (Wang et al., 2024; Sumers et al., 2023; Xi et al., 2023) are proposed to enhance the robustness and task solving abilities of LLMs for real-world complex tasks, such as web-navigation (Yao et al., 2022a; Nakano et al., 2021; Zhou et al., 2023) and software development (Hong et al., 2023; Qian et al., 2024; Yang et al., 2024). Following this promising direction, researchers design various LLM based agents for solving specific spatial-temporal tasks, including mobility prediction (Feng et al., 2024), trajectory generation (Li et al., 2024) and traffic signal control (Lai et al., 2023).
Besides, researchers also explore the potential of applying LLM based agent for automatic research experiments (Bogin et al., 2024; Romera-Paredes et al., 2024; Boiko et al., 2023) especially machine learning experiments (Huang et al., 2024; Shen et al., 2024; Zhang et al., 2024; Wu et al., 2024). For example, HuggingGPT (Shen et al., 2024) is proposed to utilize LLM to manage existing AI models in Huggingface community to solve complicated AI tasks. MLAgentBench (Huang et al., 2024) designs MLagent to automatically solve simple kaggle competition with ReAct mechanism in terms of success rate. DSBench (Zhang et al., 2024) verifies the LLM based agent for solving classic data science tasks with tabular data as input. However, trajectory data are more complicated than the tabular data or text data and different trajectory modelling tasks also exhibit diverse problem settings and models, which are too challenging for these existing agents designed for general machine learning tasks with simple data input and interface.
Recently, UrbanLLM (Jiang et al., 2024) follows the idea of HuggingGPT (Shen et al., 2024) to manage the existing specialized spatial-temporal models and APIs for tackling urban related tasks, which covers some trajectory modelling tasks. However, existing specialized spatial-temporal AI models usually designed for specific task with specific data and format, directly utilizing these models can suffer significant performance drop even can not generate any make sense results. In this paper, different from general machine learning agent and UrbanLLM for static spatial-temporal model, our proposed agentic framework is designed for unified trajectory modelling which can provide automatically model training and optimization for various trajectory data and tasks.
5. Conclusion
In this paper, we propose TrajAgent, an LLM-based agentic framework for unified trajectory modelling. Supported by UniEnv, which provides a unified data and model interface, and AutOpt for data and parameter optimization, TrajAgent can automatically identify and train the appropriate model, delivering competitive performance across a range of trajectory modelling tasks. TrajAgent introduces a new paradigm for unified trajectory modelling across diverse datasets and tasks, paving the way for advancements in trajectory modelling applications.
Appendix A Appendix
{mdframed}
[backgroundcolor=gray!20, skipabove=skipbelow=]
Parameter Optimization
User
Please select proper combination of hyperparameters of model in CONFIG HYPERPARAMETERS.Adjust the selection, the combination of hyperparameters based on the main function of hyperparameters, the characteristics of the input data, the tuning principles, and memory to get a high score. You should solve the task with interleaving Thought, Action and Observation steps.
¡CHARACTERISTICS OF INPUT DATA¿
The input temporal data contains a time dictionary(key is the user ID,the value is a list containing all time points when the user is active in chronological order) , the input user data contains a user dictionary(key is the user ID,the value is a list containing all items that the user interacts with in chronological order).
¡CONFIG HYPERPARAMETERS¿
{loc_emb_size: 500, tim_emb_size: 10, ...}
¡TUNING PRINCIPLES¿
1.Start with a small batch size (32-64) and a small learning rate (0.001-0.01): This will help prevent overshooting and overfitting.
2.Increase batch size and learning rate: If the model is not overfitting, increasing the batch size and learning rate can help improve convergence.
3.Add dropout (0.2-0.5) to prevent overfitting: If the model is overfitting, adding dropout can help regularize the model.
4.Increase embedding size: If the model is not capturing enough information, increasing the embedding size can help improve representational power.
5.Decrease learning rate and increase batch size: If the model is not converging, decreasing the learning rate and increasing the batch size can help improve stability.
¡MEMORY¿:
¡SCRATCHPAD¿:
In Thought step,you should reason how to choose the combination of hyperparameters to get a higher score.Please consider following aspects: 1.Observe the hyperparameters with high scores in MEMORY, to determine the optimal hyperparameters.
2.Use a grid search: Perform a grid search over a range of hyperparameters with high scores to find the optimal combination with higher scores.
3.Increase batch size and learning rate, and use dropout to avoid overfitting.
4. Stop or reverse the adjusting trend if the score is decreasing. According to above aspects,please first learn experiences from MEMORY, then make plan for the action step.Please use the sentence structure ’Firstly… Then… Lastly’.Let’s think step by step.
Thought:
Assistant
Firstly, I will analyze the current hyperparameters in the CONFIG HYPERPARAMETERS. …
Then, I will consider the hidden size, which is currently set at 500. …
Next, the learning rate is currently at 0.001, which is a good starting point. …
Lastly, I will monitor the model’s performance closely….
Now, I will proceed with the action step to adjust the hyperparameters based on this reasoning.
Action:
1. Increase the tim_emb_size from 10 to 20 to capture more temporal information.
2. Keep the loc_size at 500 for now, as it seems adequate.
3. Increase the hidden_size from 500 to 600 to enhance the model’s capacity.
4. Increase the dropout_p from 0.3 to 0.4 to further prevent overfitting.
5. Set the learning_rate to 0.005 to improve convergence speed.
6. Start with a batch size of 32 and monitor the performance closely.
Observation:
After implementing these changes, I will evaluate the model’s performance on a validation set. …
User
In Action step, you should consider the Thought step in SCRATCHPAD, and give a dict:{hypermeter name:hypermeter value}.The hypermeter name should be the same with the raw config hypermeter names in CONFIG HYPERPARAMETERS,and hypermeter values should be the same type as the hypermeter values in CONFIG HYPERPARAMETERS.Please do not add any comments to each value.
Action:
Assistant
{’loc_emb_size’: 500,…}
{mdframed}
[backgroundcolor=gray!20, skipabove=skipbelow=]
Task Understanding
User
The description of each task is in TASK_DESCRIPTION.
Please parse out the task name the sentence aims to address in RAW_INSTRUCTS.
1. The task name should match the key in the DESCRIPTION.
2. Please only output the task name.Do not output other contents.
¡TASK_DESCRIPTION¿
{Next_Location_Prediction: …,Trajectory_User_Linkage:…}
¡RAW_INSTRUCT¿
I want to identify the trajectory…
Assistant
Trajectory_Completion
{mdframed}
[backgroundcolor=gray!20, skipabove=skipbelow=]
Data Augmentation
User
¡TASK¿
Please:
1.select proper augmentation methods and use them in proper order to jointly augment the input temporal data and user data.Please adjust the selection and combination sequence of operators based on MEANING OF OPERATORS, CHARACTERISTICS OF INPUT DATA, and MEMORY to get a high score.
2.select proper combination of hyperparameters of each augmentation method in CONFIG HYPERPARAMETERS.Adjust the selection, the combination of hyperparameters based on the main function of hyperparameters, the characteristics of the input data, and memory to get a high score.
You should solve the task with interleaving Thought, Action and Observation steps.
¡CONFIG HYPERPARAMETERS¿
1.Ti-crop: crop_nums:(int) the size of cropping. default is 2.
crop_ratio:(float) the ratio of cropping. default is 0.
crop_n_times:(int) the number of cropping for each sequence for default setting. default is 2.
crop_time_sort:(str, choice in [maximum, minimum]) choose the candidate subsequence in a descending/ascending order according to its time interval variance. default is ’minimum’. …
2.Ti-insert_unvisited: …
¡MEANING OF OPERATORS¿
1.Ti-crop: For each user, filter out sessions with fewer trajectory points than ti_threshold + ti_crop_n_times. In each session, randomly extract several segments (ti_crop_n_times) of equal length (ti_threshold). Sample data segments according to the rules.
2.Ti-insert_unvisited:…
¡MEMORY¿:
¡SCRATCHPAD¿:
In Thought step,you should reason how to choose the combination of operators and proper combination of hyperparameters of each augmentation method to get a higher score.Please consider following aspects:
1.The meaning of each operator,and the impact of adding or removing operators at a specific location on the score.
2.Common characteristics of operator index lists with higher scores in MEMORY.Avoid using the same index list as MEMORY with scores lower than 0.18111313868613138.
3.The meaning of hyperparameters of each operator you select, and how to adjust the hyperparameters based on the characteristics of input data and hyperparameters with high scores in MEMORY.
4.Use a grid search: Perform a grid search over a range of hyperparameters with high scores to find the optimal combination with higher scores.
7. Stop or reverse the adjusting trend if the score is decreasing.
According to above aspects,please first learn experiences from MEMORY, then make plan for the action step.Please use the sentence structure ’Firstly… Then… Lastly’.Let’s think step by step.
Thought:
assistant
**Thought:**
Firstly, I will analyze the characteristics of the input data, which consists of user activity timestamps and the corresponding items they interacted with. This suggests that temporal relationships and user preferences are crucial for effective augmentation. The augmentation methods should enhance these aspects without distorting the underlying patterns.
Then, I will consider the meaning of each operator and their potential impact on the score. For instance, using ‘Ti-crop‘ can help in segmenting the data, which is essential for understanding user behavior over time. …
Lastly, I will select hyperparameters for each operator based on their functions and the characteristics of the input data.
**Action:**
1. **Select Augmentation Methods:** - **Ti-crop**:
- ‘crop_nums‘: 3
- ‘crop_ratio‘: 0.1 (to allow for a small portion of the data to be cropped)
- ‘crop_n_times‘: 3
2. **Hyperparameter Configuration:**
- For ‘Ti-crop‘, I will keep ‘crop_n_times‘ at 3 to ensure multiple segments are created….
3. **Evaluate and Adjust:**
- After implementing the above configurations, I will evaluate the score…
**Observation:**
I will monitor the score after each adjustment to ensure that the changes lead to an improvement. If the score decreases, I will revert to the previous settings.
User
¡SCRATCHPAD¿
**Action:**
1. **Select Augmentation Methods:**
- **Ti-crop**: …
In Action step, you should consider the Thought step in SCRATCHPAD, and return a list and a dictionary.
The list should contain the indices of augmentation methods in {1: ’Ti-crop’, 2: ’Ti-insert_unvisited…}.For example, if you want to first use Ti-crop, secondly use Ti-insert_random, thirdly use Ti-mask, then the list should be ’[1,4,8]’.
The dict:….The hypermeter name should be the same with the corresponding operator hypermeter names in CONFIG HYPERPARAMETERS,and hypermeter values should be the same type as the corresponding operator hypermeter values in CONFIG HYPERPARAMETERS.
Please directly output the list and dictionary.For example:…
Action:
Assistant
[1, 3, 5]
{{crop_nums:3,crop_ratio:0,….},3:{…},5:{…}}

### v2

@@
-TrajAgent: An Agent Framework for Unified Trajectory Modelling
+TrajAgent: An LLM-based Agent Framework for Automated Trajectory Modeling via Collaboration of Large and Small Models
 Abstract.
 Trajectory modeling, which includes research on trajectory data pattern mining and future prediction, has widespread applications in areas such as life services, urban transportation, and public administration.
-Numerous methods have been proposed to address specific problems within trajectory modelling.
-However, due to the heterogeneity of data and the diversity of trajectory tasks, achieving unified trajectory modelling remains an important yet challenging task.
-In this paper, we propose TrajAgent, a large language model-based agentic framework, to unify various trajectory modelling tasks.
+Numerous methods have been proposed to address specific problems within trajectory modeling.
+However, the heterogeneity of data and the diversity of trajectory tasks make effective and reliable trajectory modeling an important yet highly challenging endeavor, even for domain experts.
+In this paper, we propose TrajAgent, a agent framework powered by large language models (LLMs), designed to facilitate robust and efficient trajectory modeling through automation modeling.
+This framework leverages and optimizes diverse specialized models to address various trajectory modeling tasks across different datasets effectively.
 In TrajAgent, we first develop UniEnv, an execution environment with a unified data and model interface, to support the execution and training of various models.
-Building on UniEnv, we introduce TAgent, an agentic workflow designed for automatic trajectory modelling across various trajectory tasks.
-Specifically, we design AutOpt, a systematic optimization module within TAgent, to further improve the performance of the integrated model.
-With diverse trajectory tasks input in natural language, TrajAgent automatically generates competitive results via training and executing appropriate models.
-Extensive experiments on four tasks using four real-world datasets demonstrate the effectiveness of TrajAgent in unified trajectory modelling, achieving an average performance improvement of 15.43% over baseline methods.
+Building on UniEnv, we introduce an agentic workflow designed for automatic trajectory modeling across various trajectory tasks and data.
+Furthermore, we introduce collaborative learning schema between LLM-based agents and small speciallized models, to enhance the performance of the whole framework effectively.
+Extensive experiments on four tasks using four real-world datasets demonstrate the effectiveness of TrajAgent in automated trajectory modeling, achieving a performance improvement of 2.38%-34.96% over baseline methods.
 *Both authors contributed equally to this research.
-copyright: acmcopyrightcopyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
+copyright: acmcopyrightcopyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
 1. Introduction
-With the rapid development of web services and mobile devices (Zheng, 2015; Chen et al., 2024), large-scale trajectory data, such as check-in data from social network (Yang et al., 2016), have been collected, greatly facilitating research in trajectory modelling.
-Trajectory modelling involves the processing, mining and prediction of trajectory data, with widespread applications in urban transportation, location services and public management.
-The typical areas of trajectory modelling (Chen et al., 2024; Graser et al., 2024) can be classified into three main categories: trajectory classification (Liang et al., 2022), trajectory prediction (Yang et al., 2022), and trajectory generation (Wei et al., 2024).
+Figure 1. The paradigm of LLM based automated trajectory modeling framework TrajAgent.
+With the rapid development of web services and mobile devices (Zheng, 2015; Chen et al., 2024), large-scale trajectory data, such as check-in data from social network (Yang et al., 2016), have been collected, greatly facilitating research in trajectory modeling.
+Trajectory modeling involves the processing, mining and prediction of trajectory data, with widespread applications in urban transportation, location services and public management.
+The typical areas of trajectory modeling (Chen et al., 2024; Graser et al., 2024) can be classified into five main categories: trajectory representation (Gong et al., 2023), trajectory classification (Liang et al., 2022), trajectory prediction (Yang et al., 2022), trajectory recovery (Si et al., 2023), and trajectory generation (Wei et al., 2024).
 Each category encompasses various sub-tasks; for instance, the trajectory prediction task can be further divided into next location prediction task (Liu et al., 2016), final destination prediction task (Zhang et al., 2019), and travel time estimation task (Wang et al., 2018), among others.
-Given the huge value of trajectory modelling in diverse practical applications, various algorithms and models (Graser et al., 2024) have been proposed to address these tasks, particularly deep learning-based models in recent years.
+Given the huge value of trajectory modeling in diverse practical applications, various algorithms and models (Graser et al., 2024) have been proposed to address these tasks, particularly deep learning-based models in recent years.
 This has facilitated significant advancements in the field, with many tasks achieving a high level of performance.
@@
 For example, TrajFormer (Liang et al., 2022) is tailored for trajectory classification and cannot be applied in trajectory prediction or trajectory generation.
-Flash-back (Yang et al., 2020) is designed for sparse check-in trajectory prediction and is not suitable for dense GPS trajectory or road network-based trajectory modelling.
+Flash-back (Yang et al., 2020) is designed for sparse check-in trajectory prediction and is not suitable for dense GPS trajectory or road network-based trajectory modeling.
 In other words, due to the heterogeneity of application scenario and the diverse nature of trajectory data–varying in resolution, format and geographical regions–existing methods can only be applied in limited task with specific data for specific regions.
-Without a unified model, the deployment and application of these advanced models become costly and inefficient, hindering further development of trajectory modelling area.
-Therefore, developing a unified model for various trajectory modelling tasks is both emergent and fraught with significant challenges.
-In last two years, the rapid development of large language models (LLMs) (OpenAI, 2024; Touvron et al., 2023) with extensive commonsense and powerful reasoning abilities presents great potential for solving challenging tasks, such as mathematics (Romera-Paredes et al., 2024) and logic games (Yao et al., 2022b).
-Furthermore, LLM based agent (Xi et al., 2023; Sumers et al., 2023) have emerged as a new paradigm, offering more robust and intelligent task-solving capabilities.
-These agents have demonstrated stable and promising performance in tackling complex real-world tasks, such as web-navigation (Yao et al., 2022a; Nakano et al., 2021; Zhou et al., 2023) and software development (Hong et al., 2023; Qian et al., 2024; Yang et al., 2024).
-Leveraging the powerful understanding and reasoning abilities of LLM based agents, researchers explore the potential of utilizing LLM based agent for the unified modelling of diverse machine learning tasks (Huang et al., 2024; Shen et al., 2024; Zhang et al., 2024; Wu et al., 2024).
-For example, HuggingGPT (Shen et al., 2024) utilize LLM as a core manager to address various machine learning tasks with existing AI models, VisionLLM (Wu et al., 2024) investigates unified modelling for vision tasks across different domains, while Mora (Yuan et al., 2024) designs a multi-agent framework for replicating Sora’s functionalities with combining existing open source models.
-Inspired by these works, we explore the potential of addressing the unified trajectory modeling problem using an LLM-based agent framework.
-However, designing an LLM-based agent for unified trajectory modelling is challenging due to several factors.
-Firstly, trajectory data and existing models for various trajectory modelling tasks are diverse and often inconsistent, making the integration and deployment of these models difficult.
-Secondly, the processing steps from raw trajectory data to final output results are lengthy and cumbersome (Chen et al., 2024), resulting in a large action space for the planning and execution of entire process.
-Third, the performance of existing models highly depends one data adaptation and parameter optimization, making the direct usage of existing pre-trained models infeasible.
+While some early studies have explored effective trajectory modeling via unified framework (Lin et al., 2024; Zhu et al., 2024; Wu et al., 2024a), they often face several limitations: 1) their performance on individual tasks lags behind that of specialized models; 2) the range of supported trajectory modeling tasks remains limited; and 3) their training and inference processes are complex and non-trivial.
+Thus, despite their significant contributions, there is still a long way to go, necessitating further exploration of more effective and reliable trajectory modeling frameworks.
+In recent years, the rapid development of large language models (LLMs) (OpenAI, 2024; Touvron et al., 2023) with extensive commonsense and powerful reasoning abilities presents enable the LLM based agent (Xi et al., 2023; Sumers et al., 2023) as a new paradigm for solving complex task, such as automated software development (Hong et al., 2023; Qian et al., 2024; Yang et al., 2024) and automated machine learning tasks (Huang et al., 2024; Shen et al., 2024; Zhang et al., 2024; Wu et al., 2024b).
+For example, HuggingGPT (Shen et al., 2024) utilizes LLM as a core manager to address various machine learning tasks with existing AI models, VisionLLM (Wu et al., 2024b) investigates unified modeling for vision tasks across different vision domains.
+Inspired by these, we explore the potential of leveraging an LLM-based agent framework for automated trajectory modeling, paving the way toward effective and reliable trajectory modeling solution.
+Specifically, our approach seeks to harness the capabilities of LLMs to establish a collaborative framework between LLMs and various specialized models, enabling the automated and unified trajectory modeling.
+However, designing such an LLM-based agent for this purpose presents several significant challenges.
+Firstly, how to handle and integrate the diverse trajectory data and specialized models for different trajectory modeling tasks into a single, unified framework is non-trivial.
+Secondly, the numerous steps involved in transforming raw trajectory data into the final model output are lengthy and cumbersome (Chen et al., 2024), making full automation of the process difficult and leading to a large action space for both planning and execution.
+Finally, while model performance heavily depends on delicate and specialized data adaptation and model optimization, the ultimate challenge lies in effectively automating the optimization of these adaptation processes.
+In this paper, we propose TrajAgent, a systematic agent framework for automated trajectory modeling across diverse tasks and data.
+First, we design a unified environment, UniEnv, to process diverse trajectory data and provide a cohesive runtime environment for various trajectory modeling tasks within TrajAgent.
+In UniEnv, we define unified data and model interfaces to facilitate the seamless execution of different trajectory modeling methods.
+Building upon this environment, we develop an agentic workflow within TrajAgent for the automatic multi-step planning and execution of trajectory modeling tasks.
+The diverse trajectory modelling task workflow is decomposed into four unified steps: task understanding, task planning, task execution, and task summarization.
+For each step, we design an expert agent to perform the corresponding operations effectively.
+Finally, we introduce a collaborative learning schema that integrates agent learning through reasoning with model learning through training, enabling the effective optimization of model performance for specific data and tasks.
+In summary, our contributions are as follows,
+  • To the best of our knowledge, TrajAgent is the first LLM-based agent framework for automated and unified trajectory modeling across diverse data and tasks.
+It decomposes the trajectory modeling process into several sub-tasks, with expert agents designed to each.
+  • To support the automated execution of various trajectory modeling tasks, we provide an unified running environment UniEnv by integrating diverse trajectory data and specialized models.
+  • Furthermore, we develop a collaborative learning schema between high-level agents and low-level models to effectively enhance the final performance of proposed framework on targeted data and tasks.
+  • Extensive experiments on four representative trajectory modeling tasks across four real-world datasets demonstrate the effectiveness of the proposed framework, with the optimized model achieving a performance gain of 2.38% to 34.96% over baseline methods.
+2. Methods
 Figure 2.
-Framework of TrajAgent, including UniEnv with unified data and model interface for running experiment, TAgent an agentic workflow for executing the whole trajectory modeling procedure, and AutOpt for the model enhancement and adaptation on selected data.
-In this paper, we propose TrajAgent, a systematic agent framework for unified trajectory modelling and analytics.
-Within TrajAgent, we first design a unified environment, UniEnv, for processing diverse trajectory data and preparing a cohesive running environment for various trajectory modelling tasks.
-In UniEnv, we define unified data and model interfaces to support the easy execution of various methods for trajectory modelling.
-Next, we develop an agentic workflow called TAgent for the automatic planning and execution of trajectory modelling tasks.
-In TAgent, we decompose the task workflow into four steps: task understanding, task planning, task execution and task summarization.
-In each step, we design a specialized expert agent to perform specific operations.
-Notably, we introduce AutOpt to enhance the task execution module, enabling the adaption of model performance through various automatic optimization mechanisms, including data augmentation, parameter optimization and joint optimization.
-In summary, our contributions are as follows,
-  • To the best of our knowledge, TrajAgent is the first attempt to employ an LLM-based agentic framework for unified trajectory modeling across diverse trajectory data and various trajectory tasks.
-  • In TrajAgent, we design a agentic workflow to decompose the whole trajectory modelling task into several sub-tasks, all carried out within a unified environment UniEnv.
-Finally, we develop a comprehensive model optimization module to further enhance the model performance on targeted data.
-  • Extensive experiments on four representative trajectory modelling tasks across four real-world datasets demonstrate the effectiveness of proposed framework, which the optimized model achieving an average performance gain of 15.43%.
-2. Methods
+The framework of TrajAgent is depicted in the center, consisting of four stages: task understanding, task planning, task execution, and task summary.
+Besides, to ensure smooth operation across diverse data and tasks, we introduce UniEnv, an experimental platform with a unified data and model interface, shown on the left side of the figure.
+Finally, on the right side of the figure, we present the collaborative learning schema in TrajAgent, which includes high-level learning from knowledge reasoning for agents and low-level learning from data training for specialized models.
+2.1. Overview of TrajAgent
+Figure 1 presents the comparison between our work and existing works.
+Over the past few decades, researchers have developed various task-specific models for solving single tasks on limited datasets, as shown in the top-left of Figure 1.
+Recently, some works (Zhu et al., 2024; Lin et al., 2024; Wu et al., 2024a) have explored the potential of building unified models for multiple trajectory modeling tasks, as shown in the top-right of Figure 1.
+However, limitations in available data and the diversity of tasks have constrained these unified models to a narrow range of tasks.
+Additionally, these models are not easy to train and utilize, making optimization challenging and far from being simple and user-friendly.
+Thus, following the success of LLM-based agentic framework in other domains, we propose to build a unify framework to enable the automated trajectory modeling via the collaboration between LLM-based agents and smaller specialized models.
+The framework is presented in the bottom of Figure 1.
+In this framework, the LLM serves as a controller and processor, coordinating with specific smaller models to accomplish specific trajectory tasks.
+This approach enables users to effortlessly extend support for various data and models without delving into the intricate details of individual models.
+In essence, users can view the entire framework as a unified modeling platform, achieved through automated trajectory modeling across a variety of tasks and data.
+As TrajAgent A is to generate an optimized models M’, based on a user query q, optional trajectory data D, and the selected raw model M, the task definition is as follows,
+$M^{\prime}=\arg\min_{M}\mathcal{L}(M,T,q,D,A),$
+$\text{where }T=A(q),D=A(q,T),M=A(T,D).$
 Figure 2 presents the whole framework of TrajAgent.
-It contains three components: 1) UniEnv, an environment with a unified data and model interface that supports the execution and training of various models; 2) TAgent, an agentic workflow designed to automatically complete diverse trajectory modelling task.
-It accepts requirements in the form of natural language input and directly outputs the adapted models; 3) AutOpt, an additional optimization module for TAgent that enhances the performance of specific model through data and parameter adaptation.
+It contains three key components: (1) UniEnv, an environment with a unified data and model interface that supports the execution and training of various trajectory models; 2) Agentic Workflow, which is designed to automatically decompose and complete diverse trajectory modelling tasks; 3) Collaborative Learning Schema, an additional automated optimization module for enhancing the performance of specific model through the collaboration between LLM-based agents and specialized models with different learning mechanisms.
 Details of each component are presented as follows.
-2.1. UniEnv
+2.2. UniEnv: Environment for Experiments
 As shown in the left of Figure 2, UniEnv is a comprehensive and integrated environment that bridges trajectory data, tasks, and models, providing a foundational platform for trajectory modelling and analysis.
@@
 2.2.1. Task Interface
-We summarize the supported trajectory modelling tasks in UniEnv as follows.
-  • Trajectory prediction task: next location prediction and travel time estimation.
-  • Trajectory classification task: trajectory user linking and trajectory anomaly detection.
-  • Trajectory generation task: trajectory recovery and trajectory generation.
+Figure 3 summarizes the trajectory modeling tasks and associated models supported by UniEnv.
+The framework covers 5 fundamental trajectory modeling tasks: prediction, recovery, classification, generation, and representation, utilizing a total of 18 methods.
+For tasks such as prediction, recovery, and classification, we further introduce several subtasks.
+For example, the prediction task includes subtasks like next location prediction and travel time estimation, the classification task includes subtasks like trajectory user linking, intention prediction and anomaly detection.
 To enhance the clarity and effectiveness of understanding users’ language queries, we provide a detailed language description for each task.
@@
 2.2.3. Model Interface
-To support various model training in TrajAgent, we select at least one well-known model for each task and adapt them to match the running environments in UniEnv.
+As previously mentioned, we support 18 models across 5 fundamental trajectory modeling tasks.
+To support training these models in TrajAgent, we select at least one well-known model for each task and adapt them to match the running environments in UniEnv.
 Furthermore, we extract the semantic context from original paper of each model with the txyz.ai APIs  [^1] https://www.txyz.ai/ to generate detailed description of model.
@@
 Here, we also regard the LLM APIs utilized in the agentic workflow as one of the interface in the UniEnv, including the ChatGPT API  [^5] https://openai.com/ and DeepInfra  [^6] https://deepinfra.com for open-source LLMs.
-2.2. TAgent
-As shown in the middle of Figure 1, the architecture of TAgent is organized into four key modules: Task Understanding, Task Planning, Task Optimization, and Task Summary, which form an automated processing chain from user query to final result, eliminating the need for human-in-the-loop.
+Figure 3.
+The TrajAgent framework supports 5 fundamental trajectory modeling tasks, encompassing a total of 18 methods.
+These tasks include trajectory prediction, recovery, classification, generation, and representation.
+For the first three tasks, several common subtasks are also provided.
+Each task is equipped with at least one method, ensuring comprehensive coverage of foundational trajectory modeling objectives and SOTA algorithms.
+Detailed introduction of methods can refer to the appendix A.
+2.3. Agentic Workflow of TrajAgent
+As shown in the middle of Figure 2, the agentic workflow of TrajAgent is organized into four key modules: task understanding, task planning, task optimization, and task summary, which form an automated processing chain from user query to final result, eliminating the need for human-in-the-loop.
 Specifically, the task understanding module first receives user instructions in natural language form, and analyzes and identifies the type, name, and other key information of the tasks involved.
 Then, the task planning module will plan for the identified task, including dataset matching and model selection.
-Next, the task execution module executes the planning task and cooperate with the additional model optimization module AutOpt to further improve the task performance from both data augmentation and parameter selection perspectives.
-Last, the task summary module generate an analytical report of the task based on historical interactions and decisions of TAgent.
-Following the common practice(Shinn et al., 2024), each module in TAgent can be regarded as a small agent, which consists of a function module for executing its core function, a memory for recording the history interaction, and a reflection module for learning practice from the memory.
+Next, the task execution module executes the planning task and cooperate with the additional performance optimization module collaborative learning schema to further improve the task performance from both the agent learning and model learning perspectives.
+Last, the task summary module generate an analytical report of the task based on historical interactions and decisions of TrajAgent.
+Following the common practice(Shinn et al., 2024), each module in TrajAgent can be regarded as a small agent, which consists of a function module for executing its core function, a memory for recording the history interaction, and a reflection module for learning practice from the memory.
 2.3.1. Task Understanding Module
-As the first module of TAgent, task understanding module is designed to interact with user and extract detailed task information to launch subsequent stages.
+As the first module of TrajAgent workflow, task understanding module is designed to interact with user and extract detailed task information to launch subsequent stages.
 Given the user query, understanding module recognize the potential task name from it with the predefined supported tasks as additional input.
@@
 Follow the task understanding module is the planning module which is designed to generate the subsequent execution plan for efficient experiments of trajectory modelling.
-The input of the planning module is the task name and description from the undertanding module, the supported data and model with brief description from UniEnv.
+The input of the planning module is the task name and description from the understanding module, the supported data and model with brief description from UniEnv.
 With the carefully designed prompt, the generated execution plan will contain the data name and model name for the given task, and also the detailed model optimization plan.
@@
 Give the execution plan, the task execution module is responsible for invoking UniEnv to execute the experiment plan.
-In addition to the previously mentioned basic execution interface, another interface of this module is to call AutOpt module to complete the model optimization according to the optimization plan.
-For both interface, task execution module will give the feedback including error information for failed cases and performance metrics for success cases.
-Detailed designs of AutOpt are introduced in the following section 2.3.
+In addition to the previously mentioned basic execution interface, another interface of this module is to call collaborative learning schema module to complete the model optimization automatically.
+For both interface, the task execution module will give the feedback including error information for failed cases and performance metrics for success cases.
 2.3.4. Task Summary Module
@@
 User can also directly utilize the optimized model from the experiments via APIs for further applications.
-2.3. AutOpt
+2.4. Collaborative Learning Schema
 Due to the geospatial heterogeneity and the diversity of trajectory data, the trajectory models usually cannot be directly transferred between data and regions.
 In other words, for various data in different region, the model needs to be trained from scratch.
-Thus, the sufficient optimization of various models becomes emergent.
-In TrajAgent, we design AutOpt module to complete this task.
-As shown in the right part of Figure 1, AutOpt provides several optimization mechanisms: data augmentation, parameter optimization, joint optimization, prompt optimization and optimization via tools.
-Following Reflexion (Shinn et al., 2024), we design optim agent for each kind of optimization mechanism.
+Thus, the sufficient optimization of various models with targeted data becomes emergent.
+In TrajAgent, we design collaborative learning schema to complete this automatic optimization and generate optimized specialized models for targeted task and data.
+As shown in the right part of Figure 2, the collaborative learning framework involves two levels of learning: high-level knowledge reasoning for agents and low-level data training for specialized models.
+The high-level agent proposes training settings for the low-level models based on its expert knowledge and experimental records.
+The low-level models are then trained using these settings, and their performance metrics are reported back to the high-level agent for further collaborative learning.
+This iterative process continues until the performance meets the predefined requirements or the maximum number of exploration epochs is reached.
+2.4.1. Agent Learning via Reasoning
+Following Reflexion (Shinn et al., 2024), we design expert optim agent for learning from the experimental records via reasoning.
 The standard optim agent utilizes the history operation and related results as the feedback to update its action in the next step.
@@
 Different optimization mechanism utilize the same optim agent with different action space and optimization tips.
-2.3.1. Data Augmentation
-Based on the standard optim agent, we introduce the specific action space for data augmentation.
-Following practice from existing works (Zhou et al., 2024; Dang et al., 2023; Yin et al., 2024), we define a fixed set with ten operators for data augmentation, e.g., insert, replace, split and so on.
+Besides, as shown in Figure 1, our framework can also utilize LLM-based agent as the specialized model to complete the trajectory modelling tasks.
+Thus, we design prompt optimization agent to optimize agent based specialized models.
+We keep all experimental records in each experiment, including the raw input trajectories, the LLM’s inference results, and the reasoning process.
+We select the two best-performing trajectories along with their corresponding inference results and reasoning process as few-shot examples, which are then added to the original prompt for a re-run of the experiment.
+This process of conducting experiments and selecting results is referred to as one iteration.
+In each iteration, the few-shot examples are updated based on the experiment results.
+2.4.2. Model Learning via Training
+By collaborating with the high-level agent, the low-level specialized models learn specific trajectory patterns tailored to the target trajectory data and tasks.
+To enhance performance, we introduce several optimization techniques, including data augmentation, parameter tuning, and joint optimization.
+Data Augmentation.
+Based on the high-level optim agent, we introduce the specific action space for low-level trajectory data augmentation.
+Following practice from existing works (Zhou et al., 2024; Dang et al., 2023; Yin et al., 2024), we define a fixed set with ten operators for trajectory data augmentation, e.g., insert, replace, split and so on.
 During the optimization, the operator set with simple description is provided to the optim agent, it can select any combination of operators with their simple parameters as the action.
-After action, optim agent obtain the feedback information, e.g., performance metrics, from UniEnv to continue update its memory and action.
-2.3.2. Parameter Optimization
+Then the specialized models are trained with the augmented trajectory data and report performance metrics.
+The optim agent obtain the feedback information, e.g., performance metrics, from UniEnv to continue update its memory and action.
+Parameter Optimization.
 The action space of parameter optimization is defined based on the parameters of model itself.
@@
 While the action space is large, this kind of action space is flexible to adapt with any models.
-2.3.3. Joint Optimization
-Furthermore, we introduce the joint optimization mechanisms in AutOpt to improve the final performance.
+Joint Optimization.
+Furthermore, we introduce the joint optimization mechanisms to further improve the performance.
 Due to the different working paradigms, the direct combination of two kinds of optimizations is unsuccessful.
@@
 3. Experiments
-In this section, we first introduce the experimental settings, including the datasets, task models, and evaluation metrics.
-Next, we present the experimental results, covering the performance across multiple trajectory tasks as well as the effectiveness of agent execution.
-Following that, we perform an ablation study to assess the impact of different components within the proposed framework.
-Subsequently, we investigate the effect of key parameters for TrajAgent execution.
-Finally, we conclude with a case study to further illustrate the practical application of our approach.
+Table 1.
+Performance of representative methods across five fundamental trajectory modeling tasks.
+For the nine subtasks, only one model is presented for each.
+‘DA’ represents data augmentation, ‘PO’ denotes parameter optimization, ‘PRO’ denotes prompt optimization, ‘JO’ indicates joint optimization, $\delta$ represents performance improvements.
+Task Trajectory Prediction Trajectory Recovery Trajecory Classification Trajectory Trajectory
+SubTask Next Loc Pre TTE Recovery Map-matching User Linking Intent Prediction Anomaly Generation Representation
+Metric Acc@5 MAE MAE $Acc_{m}$ Acc@5 $Acc_{i}$ AUC MAE Acc@5
+Dataset FSQ Porto Porto Tencent FSQ Beijing Porto Earthquake FSQ
+Models GETNext LLM-ZS MulT-TTE TrajBERT GraphMM S2TUL LIMP GM-VSAE DSTPP CACSR
+Origin 0.3720 0.3110 163.12 42.71 0.2014 0.5755 0.745 0.9892 0.4611 0.31
++DA 0.3894 0.6846 0.3369
++PO 0.3995 128.57 27.78 0.2427 0.757 0.9899 0.3584 0.3466
++PRO 0.3225 0.7627
++JO 0.4002 0.3225 128.57 27.78 0.2427 0.7802 0.7627 0.9899 0.3584 0.3472
+$\delta$ 7.58% 3.69% 21.18% 34.96% 20.71% 35.57% 2.38% 0 22.27% 12%
 3.1. Settings
-3.1.1. Dataset
-  • Foursquare (FSQ): This dataset consists of 227,428 check-ins collected in New York City from 04/12/2012 to 02/16/2013, Each check-in is associated with a timestamp, GPS coordinates and corresponding venue-category.
-  • Brightkite (BGK): This dataset contains 4,491,143 check-ins of 58,228 users collected from BrightKite website.
-  • Porto: This dataset contains 1.7 million taxi trajectories of 442 taxis running in Porto, Portugal from 01/07/2013 to 30/06/2014.
-Each trajectory corresponds to one completed trip record, with fields such as taxiID, timestamp and the sequence of GPS coordinates.
-  • Chengdu: This dataset contains GPS trajectory records of Chengdu from 01/11/2016 to 30/11/2016.
-Each record includes taxiID, timestamp, longitude and latitude, collected and released by Didi Chuxing.
-  • UserQueries: To verify the effectiveness of the whole system, we utilize self-instruct method (Wang et al., 2022) with 5 seed queries to generate 300 user queries as the experiment input.
-More detailed information for the Check-in and GPS trajectory datasets are summarized in Table 1 and Table 2, respectively.
-Note that the raw datasets are typically city-scale and data-intensive.
-Loading them all into the TrajAgent framework is costly.
-In this work, we are selecting a part of data for task model training and testing during experiments.
-Table 1. Statistics of the Check-in trajectory datasets.
-Datasets Foursquare (FSQ) Brightkite(BGK)
-Num. Users 463 272
-Num. POIs 19870 50061
-Num. Trajectories 10632 22208
-Table 2. Statistics of the GPS trajectory datasets.
-Datasets Porto Chengdu
-Sampling Rate 15s 3s
-Num. Trajectories 1,710,670 5,819,383
-Avg. Trajectory Length (m) 3522.64 2857.81
-Avg. Travel Time (s) 724.20 436.12
-Latitude Range [41.1401, 41.1859] [30.6529, 30.7277]
-Longitude Range [-8.6902, -8.5491] [104.042, 104.129]
-3.1.2. Task models
-In this work, We adopt various AI models and incorporate them into TrajAgent for solving trajectory-related tasks.
-According to the type of tasks they deal with, these models can be framed in the following categories:
-  • Models for Trajectory Prediction.
-We consider the classic method RNN (Liu et al., 2016) and well-performing approaches FPMC (Rendle et al., 2010) and DeepMove (Feng et al., 2018) for the task of next location prediction.
-  • Models for Trajectory User Linkage.
-We focus on methods for solving the task of Trajectory User Linkage under check-in trajectories, such as MainTUL (Chen et al., 2022) and DPLink (Feng et al., 2019).
-  • Models for Travel Time Estimation. We consider the mainstream model DeepTTE (Wang et al., 2018) to estimate the travel time for a given GPS trajectory.
-  • Models for Trajectory Anomaly Detection. We considered the well-performing method GMVSAE (Liu et al., 2020) for detecting anomalous trajectory.
-3.1.3. Metrics
-To evaluate the performance of all models on multiple trajectory tasks, we employ the following different metrics:
-  • Acc@5 and Hit@5: Acc@5 refers to the percentage of the first five results predicted correctly. Hit@5 measures whether at least one of the top-5 predictive results is correct.
-  • MAE and AUC: Mean Absolute Error (MAE) indicates the amount of deviation from the actual values.
-Area Under ROC Curve (AUC) measures how well the model correctly distinguishes the type of the sample.
-In our experiments, Acc@5 is used to measure the accuracy of trajectory prediction and agent execution, Hit@5 is adopted for evaluating the performance of trajectory user linkage task; MAE is employed to compute the error of travel time estimation, while the metric AUC is used to assess the performance of trajectory anomaly detection.
-3.2. Main Results
-In this section, we evaluate the effectiveness of TrajAgent from two dimensions: 1) Performance of trajectory modelling tasks accomplished through TrajAgent, and 2) Execution efficiency of each stages in TrajAgent.
-3.2.1. Performance on Trajectory-related Tasks
-Here, we focus on the performances of proposed TrajAgent framework across four types of trajectory tasks: Trajectory Prediction (TP), Trajectory User Linkage (TUL), Travel Time Estimation (TTE) and Trajectory Anomaly Detection (TAD).
-The first two tasks, TP and TUL, are based on check-in trajectory data, while TTE and TAD utilize GPS trajectory data.
-The results for these tasks are summarized in Table 3 and Table 4, respectively.
-For each task, we report the model performance under four different configurations: the original model (Origin), with Data Augmentation (+DA), with Parameter Optimization (+PO), and with joint optimization (+JO).
-This setup allows us to assess the individual and combined effects of data augmentation and parameter optimization on the task performance.
-Table 3 presents the performance comparison on check-in trajectory tasks TP and TUL.
-In the TP task on FSQ dataset, DeepMove improves from 0.3422 (Origin) to 0.4018 (+JO), a 17.42% increase.
-RNN, as a simpler model, experiences the most significant gains, with accuracy increasing by 51.36% through joint optimization.
-On the BGK dataset, DeepMove continues to perform well, achieving 0.6100 (+JO), marking a 9.52% improvement.
-Across both datasets, TrajAgent’s optimization techniques consistently enhance model performance, with the FSQ dataset showing particularly notable improvements.
-For the TUL task, MainTUL and DPLink exhibit consistent improvements under TrajAgent’s optimization strategies.
-the performance of MainTUL increases from 0.4871 to 0.6121 (+JO) in FSQ, while DPLink achieves a more moderate increase of 6.08%.
-These results demonstrate the effectiveness of the TrajAgent framework, particularly in automating check-in trajectory tasks.
-The results also highlight the utility of TrajAgent in terms of integrating data enhancement and parameter optimization strategies for better problem solving.
+Here, we introduce the experimental settings, including the datasets, task models, and evaluation metrics. More detailed information can refer to appendix A.
+Data.
+We utilize the widely used Foursquare (FSQ) (Yang et al., 2016) and Brightkite (BGK) (Cho et al., 2011) in our framework as the default check-in trajectory data.
+Porto (Kaggle, 2015) and Chengdu (Chuxing, 2018) are integrated in the framework as the default GPS trajectory data.
+Besides, we use Tencent (Liu et al., 2023) as the road network based methods to support road network based tasks and Beijing (Li et al., 2024) with human labeled intention to support the mobility intention prediction task.
+Finally, to verify the effectiveness of the whole system, we utilize self-instruct method (Wang et al., 2022) with 5 seed queries to generate 300 user queries as the experiment input.
+Models.
+As shown in Figure 3, our framework supports 18 models spanning 5 core trajectory modeling tasks, which are further categorized into 9 subtasks.
+The next location prediction task includes FPMC (Rendle et al., 2010), RNN, DeepMove (Feng et al., 2018), GETNext (Yang et al., 2022), LLM-ZS (Beneduce et al., 2024).
+The travel time estimation task comprises DeepTTE (Wang et al., 2018) and MulT-TTE (Liu et al., 2020a).
+The trajectory recovery task features TrajBERT (Si et al., 2023), while the map-matching task incorporates DeepMM (Feng et al., 2020) and GraphMM (Liu et al., 2023).
+The trajectory user linking task includes DPLink (Feng et al., 2019), MainTUL (Chen et al., 2022), and S2TUL (Deng et al., 2023).
+The mobility intention prediction task is supported by LIMP (Li et al., 2024).
+The trajectory anomaly detection task employs GM-VSAE (Liu et al., 2020b), and the trajectory generation task includes ActSTD (Yuan et al., 2022) and DSTPP (Yuan et al., 2023).
+Finally, the trajectory representation method is implemented using CASCR (Gong et al., 2023).
+Metrics.
+We adopt standard practices for each task to select appropriate metrics for evaluating the performance of our framework.
+The widely used Acc@k metric is employed for next location prediction, map-matching, trajectory user linking, mobility intention prediction, and trajectory representation tasks.
+The MAE metric is utilized for travel time estimation, trajectory recovery, and trajectory generation tasks.
+Lastly, the AUC metric is specifically defined for the trajectory anomaly detection task.
+3.2. Overall Performance and Generalization Capability of TrajAgent
+In this section, we assess the overall performance of TrajAgent across various fundamental trajectory modeling tasks, as summarized in Table 1.
+Furthermore, to demonstrate the generalization capability of TrajAgent across different datasets and models, we present detailed results on diverse trajectory data using several representative models in Table 2.
+As shown in Table 1, we select at least one representative models for 9 trajectory modelling tasks to present the effectiveness of proposed framework.
+As Table 1 shows, TrajAgent supports a variety of widely-known trajectory models and demonstrates superior performance across multiple trajectory modeling tasks and trajectory datasets.
+The output of TrajAgent consistently outperforms the original methods, achieving performance gains ranging from 2.28% to 35.57%.
+For instance, in the next-location prediction task, TrajAgent harnesses its agentic workflow and collaborative learning schema for automatic modeling and optimization, leading to significant performance improvements of various models.
+Specifically, it enhances the deep learning-based model GETNext by 7.58% and the LLM-based model LLM-ZS by 3.69%.
+Performance gain in other tasks and models are much larger, for example improvement from 20.71% to 34.96% for trajectory recovery tasks.
+We observe that the improvement in the trajectory anomaly detection task is minimal, primarily due to the strong performance of the original models on the dataset.
+As shown in Table 2, we presents a performance comparison for check-in trajectory tasks, including the next-location prediction task and the trajectory user linking task, across three models and two datasets.
+The first key observation is the consistent improvement observed across tasks, models, and datasets, which highlights the potential generalization of the proposed framework.
+For different trajectory tasks, datasets, and models, TrajAgent consistently provides transferable performance improvements, ranging from 6.08% to 35.57%.
+Additionally, we observe that the performance gap between different models (e.g., the top two models for each task) on specific tasks and datasets significantly narrows from 24.88% to 9.5% following the automatic optimization of TrajAgent, emphasizing the critical role of effective data augmentation and parameter optimization.
+In other words, with the support of TrajAgent, we can easily obtain more robust and reliable task performance with less relying on the delicate model designs, data processing and parameter selection.
 Table 2.
-Comparison of task performance on check-in trajectories for TrajAgent with different configurations.
+Comparison of task performance on check-in trajectories for TrajAgent with different configurations, which demonstrate the generalization of TrajAgent across different tasks, models and datasets.
 ‘DA’ represents Data Augmentation, ‘PO’ denotes Parameter Optimization, ‘JO’ indicates Joint Optimization.
 $\delta$ represents improvements.
-Task TP TUL
-Model RNN FPMC DeepMove MainTUL DPLink
+Task Next Location Prediction Trajectory User Linking
+Model RNN DeepMove GETNext MainTUL DPLink S2TUL
 Metrics Acc@5 Hit@5
-FSQ Origin 0.1795 0.2939 0.3422 0.4871 0.7551
-+DA 0.2667 0.2939 0.4018 0.5973 0.7551
-+PO 0.1795 0.3066 0.3422 0.5691 0.7686
-+JO 0.2717 0.3066 0.4018 0.6121 0.8010
-$\delta$ 51.36% 4.32% 17.42% 25.66% 6.08%
-BGK Origin 0.4422 0.5241 0.5570 0.5908 0.8993
-+DA 0.5416 0.5275 0.5647 0.6836 0.9613
-+PO 0.5022 0.5250 0.6041 0.6683 0.9552
-+JO 0.5470 0.5275 0.6100 0.7145 0.9622
-$\delta$ 23.70% 0.65% 9.52% 20.94% 6.99%
-Table 4 presents the performance comparison on GPS trajectory tasks, specifically focusing on TTE and TAD tasks.
-The models evaluated are DeepTTE for TTE and GMVSAE for TAD, with results shown for the original configuration (Origin) and after Joint Optimization (+JO).
-For the Proto dataset, the original model has an MAE of 8.48, which significantly improves to 5.85 after joint optimization, resulting in a 31% reduction in prediction error.
-On the Chengdu dataset, the MAE decreases from 7.23 to 5.95.
-As for TAD task, we find that the AUC scores were already high, the small but consistent improvements in both datasets suggest that TrajAgent’s joint optimization can further refine model performance, even for tasks where models initially perform well.
-Table 4. Comparison of task performance on GPS trajectories for TrajAgent with different configurations.
-Task
-TTE
-TAD
-Model DeepTTE GMVSAE
-Metrics MAE AUC
-Porto origin 8.48 0.9892
-+JO 5.85 0.9899
-$\delta$ 31.01% minor
-Chendu Origin 7.23 0.978
-+JO 5.95 0.984
-$\delta$ 17.70% 0.61%
-3.2.2. Efficiency of Agentic Workflow
-We select the trajectory prediction task as an example to demonstrate the efficiency of agentic workflow.
-Table 5. Execution efficiency of TrajAgent implemented by different LLMs.
-LLM Extraction Processing Data/Model Selection Data Augmentation. Parameter Optim. Joint Optim.
+FSQ Origin 0.1795 0.3422 0.3720 0.4871 0.7551 0.5755
++DA 0.2667 0.4018 0.3894 0.5973 0.7551 0.6846
++PO 0.1795 0.3422 0.3995 0.5691 0.7686 0.7570
++JO 0.2717 0.4018 0.4002 0.6121 0.8010 0.7802
+$\delta$ 51.36% 17.42% 7.58% 25.66% 6.08% 35.57%
+BGK Origin 0.4422 0.5570 0.5324 0.5908 0.8993 0.5802
++DA 0.5416 0.5647 0.6026 0.6836 0.9613 0.6903
++PO 0.5022 0.6041 0.6116 0.6683 0.9552 0.7137
++JO 0.5470 0.6100 0.6227 0.7145 0.9622 0.7240
+$\delta$ 23.70% 9.52% 16.96% 20.94% 6.99% 24.78%
+Table 3.
+Execution success rates and task performance at each stage of the agentic workflow of TrajAgent across different LLMs.
+The Acc@5 is obtained by evaluating the next-location prediction task, with DeepMove as the default specialized model.
+LLM Extraction Processing Data/Model Selection Data Augmentation Parameter Optim. Joint Optim.
 Succ. Succ. Succ. Acc@5 Succ. Acc@5 Succ. Acc@5 Succ. Acc@5
@@
 GPT-4o-mini 95.56% 92% 100% 98.00% 90% 0.2967 85% 0.1822 96% 0.3724
+Table 4.
+Ablation study of TrajAgent.
+‘MS’ stands for Model Selection, ‘DA’ represents Data Augmentation, ‘PO’ denotes Parameter Optimization, ‘JO’ stands for Joint Optimization.
+$\downarrow$ indicates a decrease in performance, and $\uparrow$ indicates an improvement.
+Agent Variants MS DA PO JO
+Succ. Acc Succ. Acc Succ. Acc Succ. Acc
+TrajAgent 100% 98% 98% 0.3050 89% 0.1895 85% 0.3724
+w/o Reflection 100% 95$\downarrow$ 98% 0.3028$\downarrow$ 90$\uparrow$ 0.1872$\downarrow$ 82$\downarrow$ 0.3212$\downarrow$
+w/o Memory 99$\downarrow$ 80$\downarrow$ 85$\downarrow$ 0.1707$\downarrow$ 70$\downarrow$ 0.2050$\uparrow$ 68$\downarrow$ 0.1804$\downarrow$
+3.3. Ablation Study and Parameter Analysis of the Agentic Workflow
+Here, we select the next location prediction task as an example to demonstrate the efficiency of designs of agentic workflow and the effects of various LLMs in Table 3 and Table 4.
+During the experiment,we select DeepMove as the default specialized model for next location prediction task.
 Table 3 compares the execution efficiency of TrajAgent implemented by different LLMs, across each stages in the trajectory modelling workflow.
@@
 These results show that TrajAgent’s efficiency is strongly influenced by the base LLM, with high-performing models like Qwen-72B and GPT-4o-mini significantly enhancing its capabilities.
-To fully realize TrajAgent’s potential, it is essential to improve LLM performance in data augmentation and parameter optimization, as many LLMs currently struggle in these aspects.
-3.3. Ablation Study
-Table 6.
-Ablation Study of TrajAgent.
-‘MS’ stands for Model Selection, ‘DA’ represents Data Augmentation, ‘PO’ denotes Parameter Optimization, ‘JO’ stands for Joint Optimization.
-$\downarrow$ indicates a decrease in performance, and $\uparrow$ indicates an improvement.
-Agent Variants MS DA PO JO
-Succ. Acc Succ. Acc Succ. Acc Succ. Acc
-TrajAgent 100% 98% 98% 0.3050 89% 0.1895 85% 0.3233
-w/o Reflection 100% 95$\downarrow$ 98% 0.3028$\downarrow$ 90$\uparrow$ 0.1872$\downarrow$ 82$\downarrow$ 0.3212$\downarrow$
-w/o Memory 99$\downarrow$ 80$\downarrow$ 85$\downarrow$ 0.1707$\downarrow$ 70$\downarrow$ 0.2050$\uparrow$ 68$\downarrow$ 0.1804$\downarrow$
+To fully realize TrajAgent’s potential, it is essential to improve the LLM performance in data augmentation and parameter optimization, as many LLMs currently struggle in these aspects, which is the important future direction of TrajAgent.
 To assess the effectiveness of the key components in TrajAgent, we conducted an ablation study using execution efficiency as the evaluation criterion.
@@
 Overall, the combined use of the memory and reflection mechanisms is crucial for optimizing TrajAgent’s performance.
-3.4. Parameter Study
-In this section, we analyze the effects of two important parameters in the TrajAgent framework: 1) thought step: the number of steps that agent thought before taking action, and 2) memory size: the size of memory units in TrajAgent.
-To be specific, we vary the values of these parameters to investigate their impact on agent performance, and the results are illustrated in Figure LABEL:fig:param-test(a) and (b).
-As shown in Figure LABEL:fig:param-test(a), increasing thought steps significantly enhances Joint Optimization (JO) and Data Augmentation (DA).
+Due to the importance of reflection and memory in the TrajAgent, we analyze the effects of two important related parameters: 1) thought step: the number of steps that agent thought before taking action in reflection, and 2) memory size: the size of memory units in TrajAgent.
+As shown in Figure LABEL:fig:param-test:thought, increasing thought steps significantly enhances Joint Optimization (JO) and Data Augmentation (DA).
 This indicates that giving TrajAgent more reasoning time improves its overall performance.
 However, the peak performance for Parameter Optimization (PO) and JO at around 10 steps suggests that once sufficient contextual understanding is achieved, further increasing the thought steps provides diminishing returns and may not be beneficial.
-From Figure LABEL:fig:param-test(b), we find that increasing memory size initially improves performance for JO and DA, with peak accuracy reached at a memory size of 10.
+From Figure LABEL:fig:param-test:memory, we find that increasing memory size initially improves performance for JO and DA, with peak accuracy reached at a memory size of 10.
 This suggests that having sufficient memory allows TrajAgent to store and utilize relevant information effectively.
@@
 For Parameter Optimization (PO), memory size has little effect, reflecting that this task does not rely as heavily on memory records.
-3.5. Case Study
+3.4. Comparison with Automated Methods
+Figure 5.
+Compared to Optuna and UrbanLLM, TrajAgent achieves better performance in trajectory user linking tasks with S2TUL as the specialized model.
+Here, POx represents parameter optimization performed over x steps, while DAx refers to data augmentation conducted over x steps.
+Here, we compare TrajAgent with the deep learning-based AutoML method Optuna (Akiba et al., 2019) and the LLM-based automated urban task-solving framework UrbanLLM (Jiang et al., 2024).
+Using the trajectory user linking task with the S2TUL model as an example, the results are presented in Figure 5.
+UrbanLLM (represented by the blue line) is designed to directly automate the use of existing models without further optimization, resulting in the lowest performance in Figure 5.
+Meanwhile, compared to the widely used deep learning-based AutoML method Optuna (represented by the red line), which focuses on parameter optimization, TrajAgent (depicted by the yellow line) achieves superior results with fewer trial-and-error iterations.
+Furthermore, its performance can be further enhanced through joint optimization combined with automated data augmentation, resulting in a significant performance improvement of over 11.1%.
+4. Related Work
+4.1. Trajectory Modelling and Analytics
+In recent year, trajectory modelling (Zheng, 2015; Chen et al., 2024; Graser et al., 2024) makes great progress on its core research questions, including prediction (Liu et al., 2016; Feng et al., 2018; Shen et al., 2020), classification (Liang et al., 2022; Song et al., 2018), recovery (Xia et al., 2021) and generation (Yuan et al., 2022).
+Trajectory prediction is to predict the next location or the trip destination of the trajectory.
+ST-RNN (Liu et al., 2016) is the first work for applying recurrent neural network into the trajectory prediction task.
+Following this work, lots of works have been done to solve the critical issues of trajectory prediction with deep learning, for example, DeepMove (Feng et al., 2018) for multi-granularity regularity modelling, Flash-back (Yang et al., 2020) for sparse trajectory prediction, GETNext (Yang et al., 2022) for utilizing aggregated mobility pattern.
+As for the classification problem in trajectory modelling, the most well-known tasks are trajectory mode detection (Jiang et al., 2017) and trajectory user linkage (Gao et al., 2017) task.
+Various techniques, such as siamese network (Feng et al., 2019), distillation learning (Chen et al., 2022) and transformers (Liang et al., 2022) are successfully applied to discover the pattern among trajectories to solve the classification tasks.
+Finally, due to the natural limitation of trajectory data quality, trajectory data recovery (Xi et al., 2019; Xia et al., 2021) and generation (Ouyang et al., 2018) are also important tasks in trajectory modelling which are crucial for the practical application in the real world.
+Thus, follow the the trend of generative modelling, diverse generative models including GAN (Ouyang et al., 2018), diffusion models (Yuan et al., 2022; Wei et al., 2024) are utilized to improve the modelling of human dynamics.
+While these specific methods accelerate the development of trajectory modelling from different aspects, they can only handle one type of task.
+In other words, the automated and unified model for all the trajectory modelling task is still missing due to the heterogeneity of tasks and trajectory data.
+In this paper, we propose to utilize the power of LLM and agent to build a unified model framework for diverse trajectory modelling tasks, which can automatically handle various trajectory data and trajectory modelling tasks without human intervention.
+4.2. Large Language Model and Agent
+With the advent of ChatGPT (OpenAI, 2024), large language models(LLMs) with extensive commonsense and outstanding reasoning abilities have been widely explored in many domains, such as mathematics (Yao et al., 2022b), question answering (Zhuang et al., 2023), and human-machine interaction (Touvron et al., 2023).
+Furthermore, agentic framework (Wang et al., 2024; Sumers et al., 2023; Xi et al., 2023) are proposed to enhance the robustness and task solving abilities of LLMs for real-world complex tasks, such as web-navigation (Yao et al., 2022a; Nakano et al., 2021; Zhou et al., 2023) and software development (Hong et al., 2023; Qian et al., 2024; Yang et al., 2024).
+Following this promising direction, researchers also explore the potential of applying LLM based agent for automatic research experiments (Bogin et al., 2024; Romera-Paredes et al., 2024; Boiko et al., 2023) especially machine learning experiments (Huang et al., 2024; Shen et al., 2024; Zhang et al., 2024; Wu et al., 2024b).
+For example, MLAgentBench (Huang et al., 2024) designs MLagent to automatically solve simple kaggle competition with ReAct mechanism in terms of success rate.
+DSBench (Zhang et al., 2024) verifies the LLM based agent for solving classic data science tasks with tabular data as input.
+However, trajectory data are more complicated than the tabular data or text data and different trajectory modelling tasks also exhibit diverse problem settings and models, which are too challenging for these existing agents designed for general machine learning tasks with simple data input and interface.
+Recently, UrbanLLM (Jiang et al., 2024) follows the idea of HuggingGPT (Shen et al., 2024) to manage the existing specialized spatial-temporal models and APIs for tackling urban related tasks, which covers some trajectory modelling tasks.
+However, existing specialized spatial-temporal AI models usually designed for specific task with specific data and format, directly utilizing these models can suffer significant performance drop even can not generate any make sense results.
+In this paper, different from general machine learning agent and UrbanLLM for static spatial-temporal model, our proposed agentic framework is designed for unified trajectory modelling which can provide automatically model training and optimization for various trajectory data and tasks.
+5. Conclusion
+In this paper, we propose TrajAgent, an LLM-based agentic framework for automated trajectory modeling.
+Supported by UniEnv, which provides a unified data and model interface, and collaborative learning schema for joint performance optimization, TrajAgent can automatically identify and train the appropriate model, delivering competitive performance across a range of trajectory modeling tasks.
+TrajAgent introduces a new paradigm for unified trajectory modeling across diverse datasets and tasks, paving the way for advancements in trajectory modeling applications.
+Appendix A Appendix
+A.1. Datasets
+  • Foursquare (FSQ): This dataset consists of 227,428 check-ins collected in New York City from 04/12/2012 to 02/16/2013, Each check-in is associated with a timestamp, GPS coordinates and corresponding venue-category.
+  • Brightkite (BGK): This dataset contains 4,491,143 check-ins of 58,228 users collected from BrightKite website.
+  • Porto: This dataset contains 1.7 million taxi trajectories of 442 taxis running in Porto, Portugal from 01/07/2013 to 30/06/2014.
+Each trajectory corresponds to one completed trip record, with fields such as taxiID, timestamp and the sequence of GPS coordinates.
+  • Chengdu: This dataset contains GPS trajectory records of Chengdu from 01/11/2016 to 30/11/2016.
+Each record includes taxiID, timestamp, longitude and latitude, collected and released by Didi Chuxing.
+  • Tencent (Liu et al., 2023): This dataset includes both a road network and a set of vehicle trajectories collected in northeastern Beijing.
+The road network consists of 8.5K road segments and 15K edges, and the trajectory dataset contains 64K vehicle trajectories, each sampled at 15-second intervals.
+  • Beijing (Li et al., 2024): This dataset contains check-in records collected from a popular social networking platform.
+It spans a period from late September 2019 to late November 2019.It also contains intent labels annotated by human for a small dataset.
+  • UserQueries: To verify the effectiveness of the whole system, we utilize self-instruct method (Wang et al., 2022) with 5 seed queries to generate 300 user queries as the experiment input.
+More detailed information for the Check-in and GPS trajectory datasets are summarized in Table 5 and Table 6, respectively.
+Note that the raw datasets are typically city-scale and data-intensive.
+Loading them all into the TrajAgent framework is costly.
+In this work, we are selecting a part of data for task model training and testing during experiments.
+Table 5. Statistics of the Check-in trajectory datasets.
+Datasets Foursquare (FSQ) Brightkite(BGK) Beijing
+Num. Users 463 272 1566
+Num. POIs 19870 50061 5919
+Num. Trajectories 10632 22208 744813
+Table 6. Statistics of the GPS trajectory datasets.
+Datasets Porto Chengdu Tencent
+Sampling Rate 15s 3s 15s
+Num. Traj. 1,710,670 5,819,383 10,000
+Avg. Traj. Length (m) 3522.64 2857.81 2492.01
+Avg. Travel Time (s) 724.20 436.12 13903.78
+Latitude Range [41.1401, 41.1859] [30.6529, 30.7277] [40.0224, 40.0930]
+Longitude Range [-8.6902, -8.5491] [104.042, 104.129] [116.265, 116.349]
+A.2. Models
+As introduced in Figure 3, we adopt various deep learning based and LLM based models and incorporate them into TrajAgent for solving trajectory-related tasks.
+According to the type of tasks they deal with, these models can be framed in the following categories:
+  • Next Location Prediction: RNN (Liu et al., 2016), attention-based method like DeepMove (Feng et al., 2018), well-performed graph-based method like GETNext (Yang et al., 2022) and two recent proposed LLM-based methods LLM-ZS (Beneduce et al., 2024) and LLMMove (Feng et al., 2024).
+  • Travel Time Estimation: DeepTTE (Wang et al., 2018) and MulT-TTE (Liu et al., 2020a) to estimate the travel time for a given GPS trajectory.
+  • Trajectory User Linkage: widely-used DPLink (Feng et al., 2019), MainTUL (Chen et al., 2022) and S2TUL (Deng et al., 2023) are considered.
+We modified DPLink’s training approach by using publicly available sparse trajectory datasets instead of heterogeneous mobility datasets for training.
+  • Travel Intent Prediction: LIMP (Li et al., 2024), which leverages the commonsense reasoning capabilities of LLMs for mobility intention inference.
+  • Trajectory Anomaly Detection: we consider the well-performing method GMVSAE (Liu et al., 2020b), which represent different types of normal trajectories in a continuous latent space.
+  • Trajectory Generation: ActSTD (Yuan et al., 2022), which capture the spatiotemporal dynamics underlying trajectories by leveraging neural differential equations and DSTPP (Yuan et al., 2023), which defines spatial temporal point process for trajectory generation.
+  • Trajectory Recovery: TrajBERT (Si et al., 2023), which encode trajectory as sentence and train a BERT model to get representations of trajectories.
+  • Trajectory Map Matching: DeepMM (Feng et al., 2020), which proposes a data augmentation approach for map-based trajectory data, and GraphMM (Liu et al., 2023), which leverages a graph-based framework to extract features from map-based trajectory data, are considered.
+  • Trajectory Representation: CASCSR (Gong et al., 2023), which use contrastive learning method to learn trajectory representations for downstream tasks.
+A.3. Metrics
+To evaluate the performance of all models on multiple trajectory tasks, we employ the following different metrics:
+  • Acc@5 and Hit@5: Acc@5 refers to the percentage of the first five results predicted correctly. Hit@5 measures whether at least one of the top-5 predictive results is correct.
+  • $Acc_{m}$: $Acc_{m}$ is the evaluation metric which computes the average matching degree of all trajectories.For each trajectory, its matching degree is the ratio of the number of matching road segments to the number of all road segments.
+  • $Acc_{i}$: $Acc_{i}$ is used to measure the accuracy of trajectory intention inference.
+It is the ratio of the number of matching predicted intention of each check-in to the total number of check-ins.
+  • MAE and AUC: Mean Absolute Error (MAE) indicates the amount of deviation from the actual values.
+Area Under ROC Curve (AUC) measures how well the model correctly distinguishes the type of the sample.
+  • JSD: Jensen–Shannon divergence (JSD) measures the discrepancy of distributions between the generated data and real-world data.
+Lower JSD denotes a closer match to the statistical characteristics and thus indicates a better generation result.
+In our experiments, Acc@5 is used to measure the accuracy of trajectory prediction and agent execution, it is also used to measure the downstream applications of trajectory representation.$Acc_{m}$ is designed for measuring the performance of map matching task.$Acc_{i}$ is designed for measuring the performance of travel intention prediction task.
+Hit@5 is adopted for evaluating the performance of trajectory user linkage task; MAE is employed to compute the error of travel time estimation and trajectory recovery, while the metric AUC is used to assess the performance of trajectory anomaly detection.JSD and RMSE are used to measure the prediction error of the spatiotemporal domain in trajectory generation task.
+A.4. Additional results on GPS trajectory data
+Table 7. Comparison of task performance on GPS trajectories for TrajAgent with different configurations.
+Task TTE Anomaly Recovery
+Model DeepTTE MultTTE GMVSAE TrajBERT
+Metrics MAE AUC MAE
+Porto origin 8.48 129.35 0.9892 13.6667
++JO 5.85 109.85 0.9899 8.0290
+$\delta$ 31.01% 15.08% minor 41.25%
+Chendu Origin 7.23 166.29 0.978 54.8060
++JO 5.95 128.57 0.984 29.8134
+$\delta$ 17.70% 22.68% 0.61% 54.39%
+Table 7 presents the performance comparison on GPS trajectory tasks, specifically focusing on TTE and TAD tasks.
+The models evaluated are DeepTTE for TTE and GMVSAE for TAD, with results shown for the original configuration (Origin) and after Joint Optimization (+JO).
+For the Proto dataset, the original model has an MAE of 8.48, which significantly improves to 5.85 after joint optimization, resulting in a 31% reduction in prediction error.
+On the Chengdu dataset, the MAE decreases from 7.23 to 5.95.
+As for TAD task, we find that the AUC scores were already high, the small but consistent improvements in both datasets suggest that TrajAgent’s joint optimization can further refine model performance, even for tasks where models initially perform well.
+A.5. Case Study
 In this section, we analyze some cases where the TrajAgent’s optimization performance is suboptimal.
@@
 Figure 6. A representative example case of TrajAgent.
-4. Related Work
-4.1. Trajectory Modelling and Analytics
-In recent year, trajectory modelling (Zheng, 2015; Chen et al., 2024; Graser et al., 2024) makes great progress on its core research questions, including prediction (Liu et al., 2016; Feng et al., 2018; Shen et al., 2020), classification (Liang et al., 2022; Song et al., 2018), and generation (Yuan et al., 2022).
-Trajectory prediction is to predict the next location or the trip destination of the trajectory.
-ST-RNN (Liu et al., 2016) is the first work for applying recurrent neural network into the trajectory prediction task.
-Following this work, lots of works have been done to solve the critical issues of trajectory prediction with deep learning, for example, DeepMove (Feng et al., 2018) for multi-granularity regularity modelling, Flash-back (Yang et al., 2020) for sparse trajectory prediction, GETNext (Yang et al., 2022) for utilizing aggregated mobility pattern.
-Besides, to solve the travel time prediction task, another important prediction problem for trajectory modelling, deep learning models including DeepTTE (Wang et al., 2018), TTPNet (Shen et al., 2020) and CARTE (Huang et al., 2022) are proposed and achieve promising performance in various scenarios.
-As for the classification problem in trajectory modelling, the most well-known tasks are trajectory mode detection (Jiang et al., 2017) and trajectory user linkage (Gao et al., 2017) task.
-Various techniques, such as siamese network (Feng et al., 2019), distillation learning (Chen et al., 2022) and transformers (Liang et al., 2022) are successfully applied to discover the pattern among trajectories to solve the classification tasks.
-Finally, due to the natural limitation of trajectory data quality, trajectory data recovery (Xi et al., 2019; Xia et al., 2021) and generation (Ouyang et al., 2018) are also important tasks in trajectory modelling which are crucial for the practical application in the real world.
-Thus, follow the the trend of generative modelling, diverse generative models including GAN (Ouyang et al., 2018), diffusion models (Yuan et al., 2022; Wei et al., 2024) are utilized to improve the modelling of human dynamics.
-While these specific methods accelerate the development of trajectory modelling from different aspects, they can only handle one type of task.
-In other words, the unified model for all the trajectory modelling task is still missing due to the heterogeneity of tasks and trajectory data.
-In this paper, we propose to utilize the power of LLM and agent to build a unified model framework for diverse trajectory modelling tasks, which can automatically handle various trajectory data and trajectory modelling tasks without human intervention.
-4.2. Large Language Model and Agent
-With the advent of ChatGPT (OpenAI, 2024), large language models(LLMs) with extensive commonsense and outstanding reasoning abilities have been widely explored in many domains, such as mathematics (Yao et al., 2022b), question answering (Zhuang et al., 2023), and human-machine interaction (Touvron et al., 2023).
-Furthermore, agentic framework (Wang et al., 2024; Sumers et al., 2023; Xi et al., 2023) are proposed to enhance the robustness and task solving abilities of LLMs for real-world complex tasks, such as web-navigation (Yao et al., 2022a; Nakano et al., 2021; Zhou et al., 2023) and software development (Hong et al., 2023; Qian et al., 2024; Yang et al., 2024).
-Following this promising direction, researchers design various LLM based agents for solving specific spatial-temporal tasks, including mobility prediction (Feng et al., 2024), trajectory generation (Li et al., 2024) and traffic signal control (Lai et al., 2023).
-Besides, researchers also explore the potential of applying LLM based agent for automatic research experiments (Bogin et al., 2024; Romera-Paredes et al., 2024; Boiko et al., 2023) especially machine learning experiments (Huang et al., 2024; Shen et al., 2024; Zhang et al., 2024; Wu et al., 2024).
-For example, HuggingGPT (Shen et al., 2024) is proposed to utilize LLM to manage existing AI models in Huggingface community to solve complicated AI tasks.
-MLAgentBench (Huang et al., 2024) designs MLagent to automatically solve simple kaggle competition with ReAct mechanism in terms of success rate.
-DSBench (Zhang et al., 2024) verifies the LLM based agent for solving classic data science tasks with tabular data as input.
-However, trajectory data are more complicated than the tabular data or text data and different trajectory modelling tasks also exhibit diverse problem settings and models, which are too challenging for these existing agents designed for general machine learning tasks with simple data input and interface.
-Recently, UrbanLLM (Jiang et al., 2024) follows the idea of HuggingGPT (Shen et al., 2024) to manage the existing specialized spatial-temporal models and APIs for tackling urban related tasks, which covers some trajectory modelling tasks.
-However, existing specialized spatial-temporal AI models usually designed for specific task with specific data and format, directly utilizing these models can suffer significant performance drop even can not generate any make sense results.
-In this paper, different from general machine learning agent and UrbanLLM for static spatial-temporal model, our proposed agentic framework is designed for unified trajectory modelling which can provide automatically model training and optimization for various trajectory data and tasks.
-5. Conclusion
-In this paper, we propose TrajAgent, an LLM-based agentic framework for unified trajectory modelling.
-Supported by UniEnv, which provides a unified data and model interface, and AutOpt for data and parameter optimization, TrajAgent can automatically identify and train the appropriate model, delivering competitive performance across a range of trajectory modelling tasks.
-TrajAgent introduces a new paradigm for unified trajectory modelling across diverse datasets and tasks, paving the way for advancements in trajectory modelling applications.
-Appendix A Appendix
+A.6. Prompt Examples
 {mdframed}
@@
 **Thought:**
-Firstly, I will analyze the characteristics of the input data, which consists of user activity timestamps and the corresponding items they interacted with.
-This suggests that temporal relationships and user preferences are crucial for effective augmentation.
-The augmentation methods should enhance these aspects without distorting the underlying patterns.
-Then, I will consider the meaning of each operator and their potential impact on the score.
-For instance, using ‘Ti-crop‘ can help in segmenting the data, which is essential for understanding user behavior over time.
-…
+Firstly, I will analyze the characteristics of the input data, which consists of user activity timestamps and the corresponding items they interacted with. ….
+Then, I will consider the meaning of each operator and their potential impact on the score. For instance, using ‘Ti-crop‘ can help in segmenting the data, …
 Lastly, I will select hyperparameters for each operator based on their functions and the characteristics of the input data.